### v2

# Effective-medium approach to the resonance distribution of wave scattering in a random point field

Abstract
In a previous paper [Phys. Rev. A 105, 042205 (2022)], the distribution of resonance poles in the complex plane of the wavenumber $k$ associated to the multiple scattering of a quantum particle in a random point field was numerically discovered. This distribution presented two distinctive structures: a set of peaks at small $k$ when the wavelength is larger than the interscatterer distance, and a band almost parallel to the real axis at larger $k$. In this paper, a detailed theoretical study based on wave transport theory is proposed to explain the origin of these structures and to predict their distribution in the complex $k$ plane. First, it is shown that the peaks at small $k$ can be understood using an effective wave equation for the average wavefunction over the disorder. Then, that the band at large $k$ can be described by the Bethe-Salpeter equation for the square modulus of the wavefunction, which is derived from the diagrammatic method. This study is supported by careful comparisons with numerical simulations. The largest simulations revealed the presence of filaments alike quantum scars for the wavefunction in the bulk of the disordered medium.
Resonance poles; Disordered medium; Effective medium; Quantum transport; Multiple scattering; Diffusion; Random Lorentz gas; Random point field; Point scatterers; Foldy-Lax model; Bethe-Salpeter equation
I Introduction
The propagation of waves in disordered media is a subject of crucial importance in many research areas in physics such as acoustics [1, 2, 3, 4, 5], electromagnetism [6, 7, 8, 9, 10], and matter waves [11, 12, 13, 14, 15, 16, 17, 18]. These phenomena are described by wave equations with random or statistically correlated spatial heterogeneities in their propagation parameters such as the index of refraction or the potential energy function [19, 20, 21, 22, 23, 24, 25, 26, 27]. Examples are given by multiple scattering of waves in a random field of scatterers such as atoms or molecules in gases, liquids, or amorphous solids. Disorder is known to cause special effects on wave propagation, depending on the ratio of the wavelength over the mean distance between the scatterers. Pioneering works on multiple scattering by Foldy [28] and Lax [29, *Lax1952] have shown that a disordered medium may be considered under some conditions as an effective medium characterized by spatially averaged propagation parameters such as a mean refractive index. Later on, Anderson discovered that waves may be localized as the result of multiple scattering in low-dimensional disordered media [31, 32]. Otherwise, in semiclassical regimes, wave multiple scattering leads to transport ruled by the diffusion equation or the Boltzmann kinetic equation [33, 34, 35].
Disordered media may be infinite, semi-infinite, or finite with different geometries like spheres or cubes. Much work is devoted to the study of wave propagation in infinite media or the transmission of waves through a slab, a wire, or a waveguide. Less is understood about the propagation and escape of waves from inside a finite disordered open system towards its exterior. In weakly open systems such as a cavity connected to a few waveguides and terminals, the number of output channels is limited and the methods of random-matrix theory apply [36, 37, 38, 39, 40, *Goetschy2011b, 42, 43]. However, in many circumstances of great interest, the disordered system is finite and strongly open. This is the case, in particular, for a fast quantum particle emitted by a source immersed in a gaseous detector such as a cloud chamber [44, 45]. Here, a fundamental issue is to understand how the wavefunction of the emitted particle is affected by the disorder in the instantaneous configuration of the atoms composing the detector and how long the particle will travel inside the detector before escaping to the surrounding vacuum. The characteristic times of this escape process may be determined in terms of the scattering resonances of the finite disordered medium formed by the random field of atoms composing the gaseous detector.
To simplify the huge complexity of the problem, this system can be described using the model of Foldy and Lax if the atoms of the gas can be represented by point-like scatterers in the sense that their radius is assumed to be much smaller than the particle wavelength in the $s$-wave approximation [45, 46, 47, 48]. The advantage of this model is the reduction of the large multiple-scattering problem to a linear system, which can be solved numerically using modern computational resources to obtain the wavefunctions in great details. Such models may be considered as quantum random Lorentz gases, as we did in our previous papers [46, 47], where we carried out the numerical study of wave propagation in a spherical medium containing hundreds to thousands of fixed point scatterers for different values of their density with respect to the wavelength. The differential and total cross sections of such systems were studied in Refs. [46, 48]. Moreover, the distribution of the scattering resonances has also been investigated in the complex plane of the wavenumber using the resonance potential method we introduced in Ref. [47]. In this way, we observed remarkable structures at low and high energies in the distribution of scattering resonances that were not understood in the numerical exploration carried out in our previous paper [47].
The purpose of the present paper is to show that these structures in the resonance distribution can be quantitatively explained with accuracy using an extension of the effective-medium theory from real to complex values for the wavenumber. To our knowledge, such a generalization of the effective-medium theory does not seem to be known in the literature. The effective-medium wave equation, also referred to as the Dyson equation, allows us to understand the formation of the low-energy structures previously observed in the resonance distribution. Furthermore, we can show that the diffusion approximation to the transport equation, known in this research area as the Bethe-Salpeter equation [19, 20, 21, 22, 27, 49, 50, 51, 52, *Doicu2019a2, *Doicu2019a3], can explain the mean spatial profile of the wavefunction density obtained after averaging over many random configurations of the system, as well as key features of the resonance distribution at high energies. Furthermore, we also evaluate the coherence length, resulting from the process of decoherence due to the randomness of the configuration of the scatterers in the disordered medium.
This paper is organized as follows. The Foldy-Lax model used to describe the propagation of the quantum particle in a gaseous detector is presented in Sec. II.1. An example of wavefunction obtained numerically in the Foldy-Lax model for a disk-shaped random point field is shown in Sec. II.2. The scattering resonances in the complex plane of the wavenumber are defined in Sec. II.3. In Sec. III, an effective-medium theory is derived from averaging over the random configurations of the atoms to predict the distribution of peaks in the resonance density at low wavenumber. This theory is limited to a single scattering event and must be supplanted, in the large wavenumber regime, by a transport theory for the multiple scattering. This is the topic of Sec. IV which contains several steps: First the presentation of semiclassical transport equations in Secs. IV.1, IV.2, and IV.3, and then comparisons with numerical simulations in Secs. IV.4 and IV.5. Finally, conclusions are drawn in Sec. V. This paper also contains several appendices with detailed derivations, consistency checks, and additional explanations regarding the equations of the main text.
Since the models considered in this paper are valid in a space of arbitrary dimension, the volume and surface area of the unit ball in the space $\mathbb{R}^{d}$ frequently appear. They are respectively given by
$V_{d}=\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2}+1)}\qquad\text{and}\qquad S_{d}=dV_{d}=\frac{2\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\>,$ (1)
where $\Gamma(z)$ denotes the gamma function [55]. The unit $d$-ball itself will be denoted by $\mathcal{B}_{d}$, and its external border by $\mathcal{S}_{d}$.
All the numerical results presented in this paper are computed with the program MSModel [56].
II Wave scattering in random media
II.1 The Foldy-Lax model
In order to describe the multiple collisions of the quantum particle in a disordered medium, we consider a simple but powerful model originally introduced by Foldy and Lax [28, 29, *Lax1952, 27]. In this model, the wavefunction of the particle undergoes elastic collisions without loss of energy in a random configuration of fixed point scatterers representing the atoms or the molecules of a gaseous detector. This model is very convenient especially because it can be efficiently solved numerically for large disordered systems without resorting to huge discretization lattices. In this model, the wavefunction $\psi(\boldsymbol{\mathrm{r}})$ obeys the stationary Schrödinger equation
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-U(\boldsymbol{\mathrm{r}})\right]\psi(\boldsymbol{\mathrm{r}})=0\>,$ (2)
where $k=2\pi/\lambda$ is the wavenumber and $\lambda$ the wavelength. The potential $U(\boldsymbol{\mathrm{r}})$ in Eq. (2) reads
$U(\boldsymbol{\mathrm{r}})=\sum_{i=1}^{N}u(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\>,$ (3)
where $u(\boldsymbol{\mathrm{r}})$ is the short-range potential of the scatterers that we will define in details soon.
In Eq. (3), the scatterers are located at the fixed positions $\boldsymbol{\mathrm{x}}_{1},\boldsymbol{\mathrm{x}}_{2},\ldots,\boldsymbol{\mathrm{x}}_{N}$. This random point field is characterized by the joint probability distribution $P(\boldsymbol{\mathrm{x}}_{1},\ldots,\boldsymbol{\mathrm{x}}_{N})$. In this regard, the statistical average of some observable $A$ over the random configurations of the scatterers is given by
$\begin{split}\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle&=\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{1}\cdots\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{N}\\ &A(\boldsymbol{\mathrm{r}};\boldsymbol{\mathrm{x}}_{1},\ldots,\boldsymbol{\mathrm{x}}_{N})\,P(\boldsymbol{\mathrm{x}}_{1},\ldots,\boldsymbol{\mathrm{x}}_{N})\>.\end{split}$ (4)
In particular, we define the local average density of scatterers per unit volume as
$n(\boldsymbol{\mathrm{r}})=\left\langle\sum_{i=1}^{N}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\right\rangle\>.$ (5)
[…]
$n(\boldsymbol{\mathrm{r}})=\begin{cases}N/V&\text{if}~{}\boldsymbol{\mathrm{r}}\in\mathcal{B}_{d}(R)\>,\\ 0&\text{otherwise}\>,\end{cases}$ (6)
where $V=V_{d}R^{d}$ is the volume of the random point field using the notation (1). In all numerical simulations, we will use the mean interparticle distance
$\varsigma=\left(\frac{V}{N}\right)^{\frac{1}{d}}\>,$ (7)
as the unit length.
In Eq. (3), the potential $u(\boldsymbol{\mathrm{r}})$ is supposed to have a finite spatial range $b$ much smaller than the wavelength ($kb\ll 1$), so that the point scattering theory developed in Sec. III of Ref. [46] can be applied. The short-range potential $u(\boldsymbol{\mathrm{r}})$ is actually defined such that the solution of Eq. (2) for $N=1$ reads
$\psi(\boldsymbol{\mathrm{r}})=\phi(\boldsymbol{\mathrm{r}})+F(k)G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{1})\>,$ (8)
where $\phi(\boldsymbol{\mathrm{r}})=\mathop{}\!\mathrm{e}^{\mathrm{i}\boldsymbol{\mathrm{k}}\cdot\boldsymbol{\mathrm{r}}}$ is the incident plane wave, and $F(k)$ is the point scattering amplitude. The quantity $F(k)$ is an intrinsic property of the scatterer and therefore is not affected by the presence of other scatterers. We will explain more about $F(k)$ soon.
In Eq. (8), $G(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ is the free Green function which satisfies
$\left(\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}\right)G(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})\>.$ (9)
It is worth noting that the wavenumber $k$ in Eq. (9) is a complex number ($k={k}_{\rm r}+\mathrm{i}{k}_{\rm i}$). This complex nature plays an important role in this paper. The solution of Eq. (9) reads
$G^{\pm}(k,r)=-\frac{1}{2\pi}\left(\frac{\mp\mathrm{i}k}{2\pi r}\right)^{\frac{d-2}{2}}K_{\frac{d-2}{2}}(\mp\mathrm{i}kr)\>,$ (10)
where $K_{\nu}(z)$ is the modified Bessel function [55] and $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$. The free Green function (10) is characterized by the asymptotic behavior
$G^{\pm}(k,r)\xrightarrow{r\rightarrow\infty}\pm\frac{1}{2\mathrm{i}k}\left(\frac{\mp\mathrm{i}k}{2\pi r}\right)^{\frac{d-1}{2}}\mathop{}\!\mathrm{e}^{\pm\mathrm{i}kr}\>.$ (11)
The symbol $F(k)$ in Eq. (8) stands for the point scattering amplitude [46]
$F(k)=\frac{1}{\pi\mu(k)(\mathrm{i}-\cot\delta(k))}\>.$ (12)
In Eq. (12), $\delta(k)$ is the $s$-wave phase shift [57, 58, 59, 60] and $\mu(k)$ is the local density of states per unit of $k^{2}$ in free space given by
$\mu(k)=\left\langle\boldsymbol{\mathrm{r}}\right|\delta(k^{2}-\hat{\boldsymbol{\mathrm{p}}}^{2})\left|\boldsymbol{\mathrm{r}}\right\rangle=\frac{S_{d}k^{d-2}}{2(2\pi)^{d}}\>,$ (13)
where $\hat{\boldsymbol{\mathrm{p}}}=-\mathrm{i}\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}$ is the momentum operator. The cross section of the point scatterer is related to the scattering amplitude $F(k)$ by [46]
$\sigma_{\rm pt}(k)=\frac{\pi}{k}\mu(k)\left|F(k)\right|^{2}\>.$ (14)
In all numerical simulations, unless otherwise mentioned, we choose to maximize the scattering amplitude (12) by imposing $\cot\delta(k)=0$:
$F_{\max}(k)=\frac{1}{\mathrm{i}\pi\mu(k)}\>.$ (15)
The point scattering model (15) has the particularity of saturating the upper bound for the point cross section [46]
$\sigma_{\max}(k)=\frac{1}{\pi k\mu(k)}\>.$ (16)
Equation (15) thus produces the largest cross section possible for any given value of $k$. On the one hand, this is useful to effectively exploit each individual scatterer, especially in small point fields ($N\lesssim 10^{2}$). On the other hand, this model should be used with caution at small values of $k$ because the cross section (16) diverges as $\operatorname{\mathcal{O}}(k^{1-d})$ in dimensions $d\geq 2$ in a non-physical way. Another feature of the model (15) is the absence of resonance corresponding to the single-point scattering.
We have now all the ingredients to solve the full wave equation (2). Given the form of the potential (3), the particle wavefunction $\psi(\boldsymbol{\mathrm{r}})$ should read
$\psi(\boldsymbol{\mathrm{r}})=\phi(\boldsymbol{\mathrm{r}})+\sum_{i=1}^{N}a_{i}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{i})\>,$ (17)
The wave amplitudes $a_{i}~{}\forall i\in\{1,\ldots,N\}$ in Eq. (17) satisfy the self-consistent equation
$a_{i}=F(k)\left(\phi(\boldsymbol{\mathrm{x}}_{i})+\sum_{j(\neq i)}^{N}a_{j}G^{+}(k,\boldsymbol{\mathrm{x}}_{i}\mid\boldsymbol{\mathrm{x}}_{j})\right)\>.$ (18)
Equation (18) can be rewritten in matrix form using the vector notations $\boldsymbol{\mathrm{\phi}}={\left(\phi(\boldsymbol{\mathrm{x}}_{1}),\phi(\boldsymbol{\mathrm{x}}_{2}),\ldots,\phi(\boldsymbol{\mathrm{x}}_{N})\right)}^{\intercal}$ and $\boldsymbol{\mathrm{a}}={\left(a_{1},a_{2},\ldots,a_{N}\right)}^{\intercal}$. The result is
$\mathsf{M}(k)\,\boldsymbol{\mathrm{a}}=\boldsymbol{\mathrm{\phi}}\>,$ (19)
where the $N\times N$ multiple-scattering matrix $\mathsf{M}(k)$ is defined by
$\mathsf{M}(k)=F(k)^{-1}\mathsf{1}-\mathsf{G}^{+}(k)\>,$ (20)
and $\mathsf{1}$ is the identity matrix. The free Green matrix $\mathsf{G}^{+}(k)$ in Eq. (20) is given by
$G^{+}_{ij}(k)=G^{+}(k,\boldsymbol{\mathrm{x}}_{i}\mid\boldsymbol{\mathrm{x}}_{j})(1-\delta_{ij})\>.$ (21)
The wavefunction is thus obtained from Eq. (17) where the coefficients $a_{i}=[\mathsf{M}(k)^{-1}\boldsymbol{\mathrm{\phi}}]_{i}$ are given by the inversion of the matrix (20). The solution of the linear system (19) and the corresponding wavefunction (17) can be computed using the program MSModel [56].
II.2 Single-configuration wavefunction
An example of the square modulus of the wavefunction for a large random configuration of $N=10^{4}$ scatterers is shown in Fig. 1(a), using the incident wave
$\phi(\boldsymbol{\mathrm{r}})=G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\>.$ (22)
The incident wave (22) represents a quantum particle created at the center of the medium ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$).
Figure 1: Square modulus of the wavefunction computed with the program MSModel [56] in a 2D disk-shaped random point field for a single configuration of $N=10^{4}$ scatterers with $k=10\,\varsigma^{-1}$. The point source is at the center ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$) and the mean free path is $\ell_{\rm s}=2.5\,\varsigma$. The circle depicts the medium boundary. (a) 4-megapixel overview. Total computation time is about 80 minutes on an Intel Core i7 processor with four threads. (b) Enlargement of the central source point. The scatterers are depicted by small black dots. (c) Enlargement of a presumed quantum scar.
The point-like scatterers are depicted as dots in Fig. 1. As seen in this figure, the density of the single-configuration wavefunction strongly fluctuates at the scale of the wavelength. In particular, the scale of the angular fluctuations of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$ out of the medium ($\left\|\boldsymbol{\mathrm{r}}\right\|\gg R$) is predicted by partial-wave theory to be
$\Delta\theta=\frac{\pi}{kR}\>,$ (23)
as given in Eq. (137) of Ref. [46]. In the case of Fig. 1(a), this gives a fluctuation scale of the order of $\Delta\theta\simeq 0.3^{\circ}$.
The figure also shows that, within the medium, the wavefunction presents disordered meandering nodal lines, which are separated on average by a distance of the order of half the wavelength, i.e., about $\pi/k$. In addition, we observe filament-shaped enhancements of the magnitude of the wavefunction density. These structures look alike the quantum scars observed in the wavefunctions of classically chaotic systems, where the wavefunction magnitude may be enhanced by constructive interferences in the vicinity of unstable classical periodic orbits [61, *Heller2001]. There is no such orbit in the present model, since the cross section of points is classically equal to zero. However, the wave nature of the quantum particle allows a nonzero cross section, which is at most of the order of $\lambda^{d-1}$, as we have seen in Eq. (16). The cross section is thus nonzero for $\lambda\neq 0$, so that the role of unstable periodic orbits may here be played by closed paths joining several point scatterers, which might explain the similarity between the structures of nevertheless different kinds of quantum systems. It should be noted that the random point field contains [^1] There are $\textstyle\binom{N}{p}=\frac{N!}{p!(N-p)}$ ways to pick up $p$ points out of $N$. This count is then multiplied by the number $\tfrac{1}{2}(p-1)!$ of closed cycles in a complete graph of $p$ vertices.
$\mathcal{C}_{{\rm cy},p}=\begin{cases}\frac{N!}{2p(N-p)!}&\text{if}~{}p\geq 3\>,\\ \tfrac{1}{2}N(N-1)&\text{if}~{}p=2\>,\end{cases}$ (24)
closed paths involving $p$ points. The total number of closed paths thus reads
$\mathcal{C}_{{\rm cy}}=\sum_{p=2}^{N}\mathcal{C}_{{\rm cy},p}\xrightarrow{N\rightarrow\infty}\frac{(N-1)!}{2}\mathop{}\!\mathrm{e}^{1}\>.$ (25)
The number (25) is so large in the limit $N\rightarrow\infty$, that we may expect such scarring structures to occur in many possible places. It is probably the first time a phenomenon of quantum scarring is observed for such a large disordered system.
The computation of such wavefunctions can be repeated for many different random configurations of the scatterers to perform their statistical average.
II.3 Scattering resonances
This paper primarily aims at studying the distribution of resonances which are formally defined as the singular points, or poles, of the resolvent $\hat{Q}(k)$ of the complete multiple-scattering problem (2)–(3). This resolvent operator is defined by
$\hat{Q}(k)=\frac{1}{k^{2}-\hat{\boldsymbol{\mathrm{p}}}^{2}-U(\hat{\boldsymbol{\mathrm{r}}})}\quad\forall k\in\mathbb{C}\setminus\mathbb{R}\>,$ (26)
and is related to the full Green function by $Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})=\left\langle\boldsymbol{\mathrm{r}}\right|\hat{Q}(k)\left|\boldsymbol{\mathrm{r}}^{\prime}\right\rangle$. The latter function describes the propagation from $\boldsymbol{\mathrm{r}}^{\prime}$ to $\boldsymbol{\mathrm{r}}$ in presence of the random point field. The poles of Eq. (26) on the positive imaginary semiaxis of the wavenumber ($\operatorname{Im}k>0$) correspond to the bound states of the system. The resolvent (26) also displays a continuum of poles, also known as a branch cut, on the real axis of $k$ which is interpreted as the continuum of (unbound) scattering states. To investigate on the characteristics of the scattering states, in particular to study the time required for the quantum particle to escape the system, it is possible to achieve an analytic continuation through this branch cut. In general, this continuation reveals other poles in the vicinity of the real axis of $k$, but in the lower half-plane ($\operatorname{Im}k<0$), which are interpreted as resonant states. The imaginary part of these poles gives the escape rate of the particle from the system. Indeed, if we assume that the wavefunction behaves as $\psi(t)\propto\mathop{}\!\mathrm{e}^{-\mathrm{i}\omega t}$, then the square modulus will decay as $\left|\psi(t)\right|^{2}=\mathop{}\!\mathrm{e}^{2\operatorname{Im}(\omega)t}=\mathop{}\!\mathrm{e}^{-\Gamma t}$. Therefore, we identify the escape rate as
$\Gamma=-2\operatorname{Im}[\omega({k}_{\rm r}+\mathrm{i}{k}_{\rm i})]=-2v({k}_{\rm r}){k}_{\rm i}+\operatorname{\mathcal{O}}({k}_{\rm i}^{3})\>,$ (27)
where $k={k}_{\rm r}+\mathrm{i}{k}_{\rm i}$ is the position of the resonance pole, and $v(k)=\partial_{k}\omega(k)$ is the group velocity. Since ${k}_{\rm i}<0$ for a resonance pole, the escape rate (27) is positive.
The advantage of the Foldy-Lax model is to provide a systematic way of defining the complex resonance poles associated to the multiple scattering of the wave in the random point field. Indeed, using the multiple-scattering matrix (20), the full resolvent (26) reads
$\hat{Q}(k)=\hat{G}(k)+\sum_{i,j}^{N}\hat{G}(k)\left|\boldsymbol{\mathrm{x}}_{i}\right\rangle[\mathsf{M}(k)^{-1}]_{ij}\left\langle\boldsymbol{\mathrm{x}}_{j}\right|\hat{G}(k)\>.$ (28)
Due to Eq. (28), the resonance poles which interest us are the solutions of the determinantal equation
$\det\mathsf{M}(k)=0\quad\text{for}~{}k\in\mathbb{C}\>.$ (29)
Equation (29) has infinitely many roots in the complex plane of $k$, and the density $\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})$ of these roots increases with the number $N$ of scatterers. The average of this two-dimensional distribution over the random configurations of the disorder can be obtained with the resonance potential method introduced in Ref. [47]. This average resonance distribution displays the structures seen, for instance, in Fig. 6 of Ref. [47].
III Effective-medium approach
In this section, we develop a theory for the low-energy peaks in the complex resonance density observed in our previous paper [47]. These peaks appear at low energy, that is in a regime of large wavelength compared to the typical interscatterer distance ($k\varsigma\ll 1$). In this regime, we expect that the wave is no longer able to resolve the spatial heterogeneities of the medium, and thus perceives a uniform effective medium [26]. A possible approach to derive an effective-medium wave equation is to average directly the wavefunction over the random configurations of the scatterers, hence leading to a Dyson-type wave equation [19, 20, 21, 26, 25, 52, *Doicu2019a2, *Doicu2019a3]. As we will see, this equation successfully predicts the locations of the resonance peaks at low energy, providing a quantitative understanding of the numerical observations in our previous paper [47].
III.1 Effective-medium wave equation
The idea of the effective-medium approach is to define the amplitude function
$A(\boldsymbol{\mathrm{r}})=\sum_{i=1}^{N}a_{i}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\>,$ (30)
which will hopefully tend to a continuous function upon averaging over the random configurations of the scatterers. Let us express Eq. (18) in terms of $A(\boldsymbol{\mathrm{r}})$. Multiplying both sides by $\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})$ and summing over $i$ gives us
$\begin{split}A(\boldsymbol{\mathrm{r}})&=F(k)\left(\sum_{i=1}^{N}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\phi(\boldsymbol{\mathrm{r}})\right.\\ &\left.+\sum_{i=1}^{N}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{x}}_{i})\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right)\>.\end{split}$ (31)
The sum over $i$ can be averaged over the random configurations. Since we assume that the point field is uniform with density $n$ given by Eq. (5), Eq. (31) averages to
$\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle=n(\boldsymbol{\mathrm{r}})F(k)\left(\phi(\boldsymbol{\mathrm{r}})+\left\langle\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right\rangle\right)\>.$ (32)
Note that, to get Eq. (32), we neglected the statistical correlation between $a_{j}$ and $\boldsymbol{\mathrm{x}}_{i}$ in the last term of Eq. (31). In order to close the equation for $\langle A(\boldsymbol{\mathrm{r}})\rangle$, we need to achieve another approximation on the last term of Eq. (32). We will assume that, under the configurational average, the sum over $\boldsymbol{\mathrm{x}}_{j}$ can be replaced by an integral
$\left\langle\sum_{j(\neq i)}^{N}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}}_{j})a_{j}\right\rangle\simeq\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (33)
The integral in the right-hand side of Eq. (33) ignores the exclusion of the $i$-th scatterer in the sum of the left-hand side. To some extent, this difference could be compensated by a factor $(N-1)/N$ in the right-hand side, but we deliberately omit it here because the number of scatterers is supposed to be large enough to make this factor tend to 1 [28, 29]. Finally, substituting Eq. (33) into Eq. (32) leads to
$\begin{split}\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle&=n(\boldsymbol{\mathrm{r}})F(k)\\ &\times\left(\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\right)\>.\end{split}$ (34)
This is a self-consistent equation for $\langle A(\boldsymbol{\mathrm{r}})\rangle$. However, we want an equation for the average wavefunction itself. We return for this purpose to Eq. (17) and we average it over the configurations. The result is
$\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle A(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (35)
Note that, in contrast to Eq. (34), the average of Eq. (17) is not subject to any approximation. This is due to the fact that Eq. (17) refers to individual scatterer positions $\boldsymbol{\mathrm{x}}_{i}$, whereas Eq. (31) involves pairs of scatterers $(\boldsymbol{\mathrm{x}}_{i},\boldsymbol{\mathrm{x}}_{j})$. When Eq. (35) is compared to Eq. (34), one notices they are directly proportional to each other. One finds
$\left\langle A(\boldsymbol{\mathrm{r}})\right\rangle=n(\boldsymbol{\mathrm{r}})F(k)\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle\>.$ (36)
Therefore, Eqs. (34) and (36) imply that the average wavefunction obeys the self-consistent equation
$\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=\phi(\boldsymbol{\mathrm{r}})+\int_{\mathbb{R}^{d}}G^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})F(k)\left\langle\psi(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (37)
It is worth noting that this equation closely resembles the general Lippmann-Schwinger equation but with $n(\boldsymbol{\mathrm{r}})F(k)$ in the role of the potential. Therefore, this means that the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ is effectively governed by the same Schrödinger equation (2) as the actual wavefunction $\psi(\boldsymbol{\mathrm{r}})$, but with $U(\boldsymbol{\mathrm{r}})$ replaced by $n(\boldsymbol{\mathrm{r}})F(k)$. From Eq. (37), one deduces the Dyson-type effective-medium equation [19, 20, 21, 26, 25, 52, *Doicu2019a2, *Doicu2019a3]
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-n(\boldsymbol{\mathrm{r}})F(k)\right]\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=0\>.$ (38)
However, the important difference between $U(\boldsymbol{\mathrm{r}})$ and $n(\boldsymbol{\mathrm{r}})F(k)$ is that the latter is complex while the former is real. As a consequence, Eq. (38) possesses exponentially evanescent solutions. Of course, this does not mean that the particle is absorbed in the medium, since the probability is strictly conserved in the model. In fact, this exponential evanescence of the wave is due to the destructive interferences that results from the disorder. The non-vanishing part of the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ can be interpreted as the coherent wave, because it keeps in phase with the incident wave $\phi(\boldsymbol{\mathrm{r}})$ [29, 63, 26, 27, 52, *Doicu2019a2, *Doicu2019a3].
III.2 Disorder-averaged Green function
In wave transport theory, a more important quantity than the average wavefunction $\langle\psi(\boldsymbol{\mathrm{r}})\rangle$ is the full Green function averaged over the realizations of the disorder. As shown in Appendix A using the diagrammatic method, the average Green function $\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ obeys the effective-medium equation (38) but with a point source at $\boldsymbol{\mathrm{r}}^{\prime}$ in the right-hand side:
$\left[\nabla^{2}_{\boldsymbol{\mathrm{r}}}+k^{2}-n(\boldsymbol{\mathrm{r}})F(k)\right]\left\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})\>.$ (39)
Regarding the analytic properties of the average Green function at complex values of $k$, one notices the similarity between Eq. (39) and Eq. (9) for the free Green function $G(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$. Accordingly, Eq. (39) has two solutions valid for $k\in\mathbb{C}$: $\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially vanishes for $\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|\rightarrow\infty$ on the domain $\operatorname{Im}k>0$, and $\langle Q^{-}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$ which exponentially increases on the same domain.
Several additional comments should be made about the average Green function $\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\rangle$. First of all, if the medium is uniform and extends to infinity, then the average Green function is essentially the same as the free-space Green function
$\left\langle Q^{+}(k,r)\right\rangle=G^{+}(\kappa(k),r)\>,$ (40)
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$, but with $k$ replaced by the effective wavenumber [28, 29, 63, 26, 25]
$\kappa(k)=\sqrt{k^{2}-nF(k)}\>.$ (41)
Due to the complex nature of the point scattering amplitude $F(k)$, the effective wavenumber $\kappa$ is itself complex, meaning that the average Green function exponentially decays with the distance $r$ from the point source.
Figure 2: Real part of the Green function in a 2D disk-shaped random point field for $N=300$ and $k=6\,\varsigma^{-1}$ with a source at the center ($\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{0}}$). The solid curve is the numerical average over $2^{14}$ configurations, the dashed curve is the theoretical average (40), and the dotted curve is the free-space Green function (10) for comparison. The filled region is the numerically computed interquartile range of $\operatorname{Re}Q^{+}(k,r)$. The radius of the medium is indicated by the gray vertical line. The point scattering model (15) is used and the mean free path is $\ell_{\rm s}=1.5\,\varsigma$. Note that $\operatorname{Im}k\xrightarrow{>}0$ for all curves, but $\operatorname{Im}\kappa(k)>0$.
This decay is numerically confirmed with the program MSModel [56] in Fig. 2 for a two-dimensional (2D) point field. The agreement between the numerical average of $\operatorname{Re}Q^{+}(k,r)$ (solid) and the theoretical average $G^{+}(\kappa(k),r)$ (dashed) is so precise that the curves are completely superimposed. The exponential decay of $\langle Q^{+}(k,r)\rangle$ at large $r$ is evidenced by the comparison with the free-space Green function $G^{+}(k,r)$ (dotted). This supports the validity of the effective-medium equation (39).
Another important property is that the characteristic length of the exponential decay of the average Green function can be physically interpreted. If we take the square modulus of Eq. (40) using the asymptotic behavior (11), we get
$\left|\left\langle Q^{+}(k,r)\right\rangle\right|^{2}\xrightarrow{r\rightarrow\infty}\frac{\pi\mu(\left|\kappa\right|)}{\left|\kappa\right|S_{d}r^{d-1}}\mathop{}\!\mathrm{e}^{-2\operatorname{Im}\kappa r}\>.$ (42)
The imaginary part of $\kappa$, given by Eq. (41), can be expanded if we assume that $k^{2}$ is relatively large compared to $nF(k)$. We find
$\begin{split}2\operatorname{Im}\kappa(k)&=2\operatorname{Im}\left(k-\frac{n}{2k}F(k)+\operatorname{\mathcal{O}}\!\left[\tfrac{1}{k^{3}}(nF)^{2}\right]\right)\\ &\simeq 2{k}_{\rm i}-\frac{n}{k}\operatorname{Im}F(k)\>.\end{split}$ (43)
Due to the optical theorem, given for instance by Eq. (44) of Ref. [46], it turns out that the last term of Eq. (43) is just the inverse of the scattering mean free path [25, 26, 64]
$\frac{1}{\ell_{\rm s}}=n\sigma=-\frac{n}{k}\operatorname{Im}F(k)\>.$ (44)
The characteristic length $\ell_{\rm s}$ represents the mean distance between two successive collisions of the wave in the medium. This length plays a crucial role in transport theory. Moreover, since we consider complex wavenumbers in this work, we have ${k}_{\rm i}\neq 0$ in general. In anticipation of future calculations, we already introduce the notation
$\gamma=2{k}_{\rm i}\>.$ (45)
In this way, the behavior (42) of the average Green function reads
$\left|\left\langle Q^{+}(k,r)\right\rangle\right|^{2}\xrightarrow{r\rightarrow\infty}\frac{\pi\mu(\left|\kappa\right|)}{\left|\kappa\right|S_{d}r^{d-1}}\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)r}\>.$ (46)
The exponential decay of the form $\mathop{}\!\mathrm{e}^{-r/\ell_{\rm s}}$ in Eq. (46) is strongly reminiscent of the Beer-Lambert law for the attenuation (or extinction) of light in materials [65, 27]. Note, however, that the present decay actually corresponds to the loss of phase coherence of the wave [26, 25, 64, 52, *Doicu2019a2, *Doicu2019a3], because, due to the elastic collisions with the scatterers, the wave looses memory of its phase and thus of its direction of motion. As mentioned earlier, this decay does not indicate the presence of absorption of the wave by the medium.
It is worth noting that the mean free path $\ell_{\rm s}=(n\sigma)^{-1}$ is longer than the mean interscatterer distance $\varsigma=n^{-1/d}$. This results from the assumption that the point field is dilute, namely the size of the scatterers, which is determined by the scattering length $\alpha$ (defined as in Refs. [46, 47]), should be smaller than the interscatterer distance ($\alpha\ll\varsigma$). An important consequence of $\ell_{\rm s}\gg\varsigma$ is that the wave is likely to pass through many scatterers before a collision effectively occurs. This is not trivial because one could a priori think of the propagation of the wave as being simultaneously affected by all the scatterers. This supposition would lead to a characteristic distance of $\varsigma$ between successive interactions, or in other words $\ell_{\rm s}\simeq\varsigma$, which turns out to be incorrect, as pointed out in Ref. [25, p. 86].
Moreover, the presence of exponential evanescence in the Green function has an important consequence concerning the random and unpredictable character of the tracks formed by a quantum particle moving in a gaseous detector [44, 45]. Indeed, Eq. (46) strongly recalls the exponential probability density
$p_{\rm col}(x)\mathop{}\!\mathrm{d}x=\frac{1}{\ell_{\rm s}}\mathop{}\!\mathrm{e}^{-x/\ell_{\rm s}}\mathop{}\!\mathrm{d}x\>,$ (47)
that the next collision takes place at the distance $x$ from the previous collision point. Therefore, each time the wave encounters a scatterer, it gets a certain probability
$P_{\rm col}=1-\mathop{}\!\mathrm{e}^{-\varsigma/\ell_{\rm s}}\simeq\frac{\varsigma}{\ell_{\rm s}}\qquad\text{for}~{}\varsigma\ll\ell_{\rm s}\>,$ (48)
to collide with it. The probability (48) never reaches one, except of course in the nonphysical limit $\ell_{\rm s}=0$. This shows that, even in quantum mechanics, the collisions constitute a Poisson process. This is an important observation because it supports the idea that the visible track of a quantum particle, such as an electron, in a gaseous detector is random and unpredictable.
III.3 Low-energy peaks in the resonance density
An important application of the effective-medium equation (39) is to provide an explanation for the low-energy peaks in the resonance density discovered in our previous paper [47]. As discussed in Sec. II.3, the resonances appear as poles of the full Green function $Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ in the lower half-plane of the wavenumber ($\operatorname{Im}k<0$). In order to facilitate the comparison between theory and simulation, we restrict ourselves to the case of a spherical-shaped disordered medium of radius $R$, and we solve the effective-medium equation (39) for the resonance poles. This calculation is shown in details in Appendix B using the partial wave method [57, 58, 59, 60]. The result is the equation
$\kappa\frac{J_{\nu+1}(\kappa R)}{J_{\nu}(\kappa R)}=k\frac{H^{+}_{\nu+1}(kR)}{H^{+}_{\nu}(kR)}\>,$ (49)
for a complex resonance at $k\in\mathbb{C}$. In Eq. (49), $\kappa$ is the effective wavenumber (41), $J_{\nu}(z)$ and $H^{+}_{\nu}(z)$ are the Bessel and Hankel functions, respectively, and the order $\nu$ is given by
$\nu=\ell+\frac{d-2}{2}\>,$ (50)
where $\ell$ is the orbital quantum number.
The comparison between the numerical solutions of Eq. (49) for $k\in\mathbb{C}$ and the actual peaks of the resonance density obtained with the program MSModel [56] is shown in Fig. 3(a) for a 2D point field and in Fig. 3(b) for a three-dimensional (3D) point field.
Figure 3: (a) Resonance density $2\pi\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})$ computed with the program MSModel [56] in a 2D disk-shaped random point field for $N=50$ and using the point scattering model given by Eq. (18) of Ref. [47] with $\alpha=0.1\,\varsigma$. The density is averaged over $2^{8}$ random configurations. The resonances of the effective medium, obtained by numerically solving Eq. (49), are depicted by small circles labelled with the corresponding angular momentum $\ell$. The dashed line separates the low from the high energy region (52). (b) Same as panel (a) but in a 3D spherical random point field for $N=100$ averaged over $2^{8}$ configurations.
In these two panels, instead of Eq. (15), we exceptionally use the point scattering model given by Eq. (18) of Ref. [47] with the scattering length $\alpha=0.1\,\varsigma$. Indeed, this point scattering model is more appropriate than Eq. (15) to explore the small-$k$ regime because the cross section does not strongly diverge for $k\rightarrow 0$.
The adequacy between the roots of Eq. (49) and the peaks in Figs. 3(a) and 3(b) is remarkable and supports the validity of the effective-medium equation (38) at large wavelength ($k\varsigma\ll 1$). It shows that, in this low-energy regime, the incident wave mainly perceives the random point field as a semitransparent sphere without the details due to disorder.
Moreover, it turns out that the resonance structure shown in Fig. 3 is little affected by the geometric shape of the disordered medium or the scattering parameters such as the cross section of the point scatterers. Indeed, the same kind of structure can be observed for a cubic medium, instead of a spherical medium. The robustness of this resonance structure allows us to make a quite radical simplification of the resonance equation (49). Indeed, given the weak influence of the cross section of the point scatterers, it is possible to let the parameter $\kappa$ tend to infinity, leading to
$H^{+}_{\nu}(kR)=0\>.$ (51)
As explained in Appendix C, this is the equation of resonances corresponding to the scattering by a hard sphere of radius $R$. The roots of Eq. (51) are shown in the figure of Appendix C and indeed form structures very close to those in Fig. 3. This observation confirms the robustness of this resonance structure against possibly significant changes of the scattering parameters.
Nevertheless, the hard-sphere resonance equation (51) is not as accurate as Eq. (49) because it neglects the scattering within the medium. This is why the roots of Eq. (51) are not shown in Fig. 3. In contrast, the prediction of the resonance equation (49) of the effective-medium theory is more reliable.
In the 2D case [Fig. 3(a)], a slight shift between simulation and prediction can be noticed. In addition, the predictions in the upper right side of Fig. 3(a) do not match the resonance band close to the real $k$ axis. These disagreements could be due to an effect of Anderson localization because, in 2D, the point cross section is not bounded at low energy due to a logarithmic singularity at $k=0$. Therefore, the product $k\ell_{\rm s}=\tfrac{k}{n\sigma(k)}$ becomes small compared to 1, which thus favors Anderson localization in a finite 2D system [31, 66, 26, 32, 64, 15].
On another side, the agreement between the theoretical estimate (49) and the numerical simulation is much more striking in the 3D case [Fig. 3(b)]. Even better: the theoretical prediction coincides with the faint peaks closest to the real axis for $\ell=0,1,2$ in Fig. 3(b). This is not obvious because these peaks constitute the leftmost edge of the resonance band which was not expected to be described by this theory. At higher energies, however, that is when
$\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}\>,$ (52)
the prediction (49) does no longer match the position of the resonance band in the upper right side of Fig. 3(b). In order to describe this band, we need to consider a more complete transport theory which remains valid at wavelengths smaller than the mean interscatterer distance ($\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$). This is the subject of Sec. IV.
IV Semiclassical transport
The issue with the effective-medium approach is that, due to the destructive interference caused by the disorder averaging, only the first collision is taken into account. When the quantum particle propagates further in the medium, multiple scattering becomes more significant. This explain why the effective-medium equation (39) is unable to describe the resonance band in Fig. 3 in the high-energy region ($\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$). To address this issue, we need to consider a more elaborate theory for wave transport.
In this section, we introduce wave transport theory and compare its prediction of the disorder-averaged square modulus of the wavefunction, $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$, to the results of numerical simulations based on the Foldy-Lax model. In particular, we want to know how good the prediction of $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$ is compared to a typical realization of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$ for a single configuration of the disorder in the regime of large system ($N\gg 1$). To this end, we let the quantum particle start from the center of the spherical medium.
Finally, we show that the wave transport theory gives an accurate prediction of the maximum of the resonance band in the high-energy region.
Throughout this section, we assume $\hbar=2m=1$ for convenience, so that the energy is also denoted as $E=\hbar\omega=\omega$, the energy-momentum relation becomes $\omega=k^{2}$, and the group velocity reads $v=2k$.
IV.1 Time evolution and complex frequencies
Besides the wavefunction, a relevant quantity in quantum mechanics is the density matrix defined for a pure state $\left|\psi(t)\right\rangle$ by
$\hat{\rho}(t)=\left|\psi(t)\right\rangle\left\langle\psi(t)\right|\>.$ (53)
This matrix contains much more information about the particle state than the probability distribution alone, especially the momentum, the direction of motion and the coherence, given by its off-diagonal elements. Moreover, the probability density is just the diagonal part of the density matrix: $\rho(\boldsymbol{\mathrm{r}},t)=\left\langle\boldsymbol{\mathrm{r}}\right|\hat{\rho}(t)\left|\boldsymbol{\mathrm{r}}\right\rangle$. The time evolution of the density matrix is governed by the quantum Liouville equation
$\mathrm{i}\frac{\partial\hat{\rho}(t)}{\partial t}=[\hat{H},\hat{\rho}(t)]\>,$ (54)
where $\hat{H}=\hat{\boldsymbol{\mathrm{p}}}^{2}+U(\hat{\boldsymbol{\mathrm{r}}})$ is the Hamiltonian of the multiple scattering problem. The initial condition of Eq. (54) at $t=0$ is $\hat{\rho}_{0}$. In the following sections, we will typically assume that the initial state is the position eigenstate
$\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|\>,$ (55)
but the present calculation is much more general. Since the Hamiltonian is time-independent, the solution of Eq. (54) simply reads
$\hat{\rho}(t)=\mathop{}\!\mathrm{e}^{-\mathrm{i}\hat{H}t}\hat{\rho}_{0}\mathop{}\!\mathrm{e}^{\mathrm{i}\hat{H}t}\>.$ (56)
The point is that, instead of calculating $\hat{\rho}(t)$ as a time-dependent function, it turns out to be more convenient to determine the Laplace transform of $\hat{\rho}(t)$ in time
$\hat{\rho}(s)=\int_{0}^{\infty}\mathop{}\!\mathrm{e}^{-st}\hat{\rho}(t)\mathop{}\!\mathrm{d}t\>.$ (57)
Another way is to use the Fourier transform of $\hat{\rho}(t)$, as usually done in the literature [25, 26]. However, we find the Laplace transform more suitable than the Fourier transform for non-oscillating signals such as the density $\rho(\boldsymbol{\mathrm{r}},t)$. The advantage of the Laplace transform $\hat{\rho}(s)$ is that it contains the same information as the time-dependent density matrix $\hat{\rho}(t)$, while being a stationary quantity.
In order to evaluate (57), we represent the evolution operators in the frequency space
$\mathop{}\!\mathrm{e}^{\pm\mathrm{i}\hat{H}t}=\int_{\mathbb{R}\mp\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\mathop{}\!\mathrm{e}^{\pm\mathrm{i}\omega t}\hat{Q}(\omega)\>,$ (58)
where $\hat{Q}(\omega)$ is the resolvent of the full problem
$\hat{Q}(\omega)=\frac{1}{\omega-\hat{H}}\qquad\forall\omega\in\mathbb{C}\setminus\mathbb{R}^{+}\>,$ (59)
introduced in Eq. (26). The parameter $\varepsilon\xrightarrow{>}0$ in Eq. (58) is an infinitesimal shift of the integration contour in the imaginary direction. This is required for convergence due to the singularities in the resolvent (59). Substituting Eq. (56) and (58) into Eq. (57) yields
$\begin{split}\hat{\rho}(s)&=\int_{\mathbb{R}+\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega}{2\pi}\int_{\mathbb{R}-\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega^{\prime}}{2\pi}\\ &\times\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\mathop{}\!\mathrm{e}^{-(s+\mathrm{i}\omega-\mathrm{i}\omega^{\prime})t}\hat{Q}(\omega)\hat{\rho}_{0}\hat{Q}(\omega^{\prime})\>.\end{split}$ (60)
Integrating over the time $t$, and using the changes of variable
$\omega_{\rm m}=\frac{\omega+\omega^{\prime}}{2}\>,\qquad\omega_{\rm d}=\omega-\omega^{\prime}\>,$ (61)
Eq. (60) becomes
$\hat{\rho}(s)=\int_{\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\omega_{\rm m}}{2\pi}\int_{\mathbb{R}+\mathrm{i}\varepsilon}\frac{\mathop{}\!\mathrm{d}\omega_{\rm d}}{2\pi}\>\frac{\hat{Q}(\omega_{\rm m}+\tfrac{\omega_{\rm d}}{2})\hat{\rho}_{0}\hat{Q}(\omega_{\rm m}-\tfrac{\omega_{\rm d}}{2})}{s+\mathrm{i}\omega_{\rm d}}\>.$ (62)
Finally, the frequency difference, $\omega_{\rm d}$, can be integrated out with the residue theorem around the simple pole at $\omega_{\rm d}=\mathrm{i}s$. The result is
$\hat{\rho}(s)=\int_{\mathbb{R}}\frac{\mathop{}\!\mathrm{d}\omega_{\rm m}}{2\pi}\>\hat{Q}(\omega_{\rm m}+\tfrac{\mathrm{i}}{2}s)\hat{\rho}_{0}\hat{Q}(\omega_{\rm m}-\tfrac{\mathrm{i}}{2}s)\>.$ (63)
To get rid of the remaining integral, we assume that the spectrum of the state $\hat{\rho}_{0}$ is very peaked around the central frequency $\omega_{0}$. Of course, one could argue that the monochromatic assumption of the state $\hat{\rho}_{0}$ is not compatible with the initial condition that $\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|$, because obviously the mean frequency of $\hat{\rho}_{0}$ cannot equal $\omega_{0}$, i.e., $\operatorname{Tr}(\hat{\rho}_{0}\hat{H})\neq\omega_{0}$, due to the uncertainty principle. However, it is possible to overcome this issue by treating $\boldsymbol{\mathrm{r}}_{0}$ as a source point for a wave at a central frequency $\omega_{0}$ during a time window of duration $\Delta t$ such that $s\ll\tfrac{2\pi}{\Delta t}\ll\omega_{0}$, as presented for instance in Ref. [67, Sec. 4.2.1]. The result is the reduction of the density matrix (63) to
$\hat{\rho}(s)=\hat{Q}(\omega_{0}+\tfrac{\mathrm{i}}{2}s)\hat{\rho}_{0}\hat{Q}(\omega_{0}-\tfrac{\mathrm{i}}{2}s)\>.$ (64)
This is the sought expression for the Laplace transform of the density matrix. It is showed in Appendix D that the normalization of the density matrix, $\operatorname{Tr}\hat{\rho}(t)$, is conserved over time, as it should in absence of absorption. This appendix also provides the absolute normalization of $\hat{\rho}(t)$ in Eq. (64), which is not 11 11.
We notice on Eq. (64) the other advantage of the Laplace transform on the usual Fourier transform [25, 26]. Indeed, the right-hand Green function in Eq. (64) is just the Hermitian conjugate of the left-hand Green function, according to $\hat{Q}^{\dagger}(\omega)=\hat{Q}({\omega}^{*})$. Therefore, the Green function only has to be evaluated at the single complex frequency $\omega_{+}=\omega_{0}+\tfrac{\mathrm{i}}{2}s$. If we had calculated the Fourier transform, $s$ would have been imaginary, leading to two different real frequencies in Eq. (64), and thus less convenient expressions. This shows the interest of the Laplace transform in this context.
The result (64) shows that the Laplace variable $s$ can be interpreted as the imaginary part of the frequency $\omega$. It is an important result because it proves the equivalence between the evolution of the density matrix in the time domain and the complex frequency domain. In addition, Eq. (64) plays a crucial role in this paper because it relates $\hat{\rho}(s)$ to the spectrum of complex resonances given by the Green operator $\hat{Q}(\omega)$.
The density matrix (64) is likely more transparent when expressed in position space. Using the initial condition (55), the density matrix reads [25, 26]
$\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)=Q^{+}(k_{+},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})Q^{-}(k_{-},\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})\>,$ (65)
or equivalently,
$\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)=Q^{+}(k_{+},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0}){[Q^{+}(k_{+},\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})]}^{*}\>.$ (66)
The wavenumbers $k_{\pm}$ in Eqs. (65)–(66) corresponding to the frequencies $\omega_{\pm}=\omega_{0}\pm\tfrac{\mathrm{i}}{2}s$ are given by
$k_{\pm}=k_{0}\pm\mathrm{i}\frac{s}{2v}\>,$ (67)
for the group velocity $v$. This also shows that $s=\gamma v$ according to the notation (45). Appendix E presents a consistency check that, in free space, Eq. (66) can indeed be interpreted as a freely propagating wave packet in the time domain.
IV.2 Semiclassical transport equation
It is possible to obtain a closed equation for the density matrix (66) after averaging over the random configurations of the scatterers, as shown in Appendix F. This leads to an integral transport equation of the same form as the Bethe-Salpeter equation [28, 68, 19, 20, 21, 24, 25, 26, 69]
$\begin{split}\left\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime},s)\right\rangle&=\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\langle{Q^{+}(k,\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})}^{*}\rangle\\ &+\int_{\mathbb{R}^{d}}\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{x}})\rangle\langle{Q^{+}(k,\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{x}})}^{*}\rangle\\ &\qquad\times\left|F(k)\right|^{2}\left\langle\rho(\boldsymbol{\mathrm{x}},\boldsymbol{\mathrm{x}},s)\right\rangle n(\boldsymbol{\mathrm{x}})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}\>,\end{split}$ (68)
where the wavenumber has the complex value $k=k_{+}$ according to Eq. (67). Equation (68) describes the multiple scattering of the particle in the disordered medium. The last term of the right-hand side accounts for the individual collisions which may occur anywhere in the medium, hence the integral over $\boldsymbol{\mathrm{x}}$.
The main assumption behind the Bethe-Salpeter equation (68) is that the wavelength is much smaller than the mean free path [^2] Note that, since the mean free path is at least equal to the mean interatomic distance ($\ell_{\rm s}=\varsigma$), the condition (69) is necessarily satisfied for $k_{0}\varsigma\gg 1$.
$k_{0}\ell_{\rm s}\gg 1\>.$ (69)
This is indeed equivalent to $\omega_{0}\gg n\sigma v$ as long as the dispersion relation is a power law: $\omega(k)\propto k^{a}~{}(\forall a>0)$. The condition (69) is known in the literature as the weak scattering regime [26], or the weak disorder regime [25]. In this regime, coherent effects such as localization play less significant roles, especially in 2D and 3D. Note that, in the context of gaseous particle detectors, the condition (69) is usually met. In this regard, some orders of magnitude are shown in Tab. 1.
Particle $k_{0}$ $\ell_{\rm s}$ $k_{0}\ell_{\rm s}$
[…]
Therefore, we find reasonable to neglect these effects in this section.
Since only the density appears in the right-hand side of Eq. (68), this equation is closed for the diagonal part of the density matrix [28], which is nothing but the probability density $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}})\rangle=\langle\rho(\boldsymbol{\mathrm{r}})\rangle=\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$. The fact that the equation is closed for $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ results from the assumption that the scatterers are point-like in the Foldy-Lax model. As a consequence, the scattering amplitude $F(k)$ is isotropic, and represented by a scalar number, instead of a momentum-dependent operator. Thus, on each collision, the particle loses the memory of its initial direction, and can be represented on each point by an isotropic spherical wave. Therefore, the only relevant distribution to describe the particle state is the density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$. However, in the more general case of directional scattering, the equation for the density matrix $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\rangle$ could not be simplified to an equation for $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$.
In the isotropic scattering case, it is convenient to define a new integral kernel with the square modulus of the effective Green function
$K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})=\frac{\left|k\right|}{\pi\mu(\left|k\right|)}\left|\left\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\right|^{2}\>.$ (70)
The integral kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ can be interpreted as the probability density for the next collision point $\boldsymbol{\mathrm{r}}$ given the previous collision happened at $\boldsymbol{\mathrm{r}}^{\prime}$. The usefulness of the prefactor in Eq. (70) is that the far-field behavior of $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ does not explicitly depend on the wavenumber $k$. In the weak scattering regime ($\left|\kappa\right|\simeq\left|k\right|$), one finds from Eq. (46)
$K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\xrightarrow{r\rightarrow\infty}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\>,$ (71)
where $\gamma=\tfrac{s}{v}=2\operatorname{Im}k$ is the dual variable of time under Laplace transform.
Using the kernel (70) and the point cross section formula (14), the diagonal part of the Bethe-Salpeter equation (68) reads
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+\int_{\mathbb{R}^{d}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (72)
Note that the small imaginary part of $k=k_{0}+\mathrm{i}{k}_{\rm i}$ can be neglected in the cross section because the equation only considers time scales much longer than the oscillation period of the wave ($s\ll\omega_{0}$ and ${k}_{\rm i}\ll k_{0}$). It is thus safe to write $\sigma(k)\simeq\sigma(k_{0})$. Note also that, from Eq. (68) to Eq. (72), we have changed the normalization of the density because the source term $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$ became $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ defined in Eq. (70). However, this has no consequence on $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ except for an irrelevant global factor.
The formal solution of the transport equation (72) reads
$\left\langle\rho(\hat{\boldsymbol{\mathrm{r}}})\right\rangle=\frac{1}{1-\hat{K}n(\hat{\boldsymbol{\mathrm{r}}})\sigma(k)}K(\hat{\boldsymbol{\mathrm{r}}}\mid\boldsymbol{\mathrm{r}}_{0})\>.$ (73)
In Sec. IV.5, we will study the spectrum of the solution (73) in the variable $\gamma$ introduced in Eq. (45), so as to make the connection with the band of complex resonances at $\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$. Indeed, according to the relationship (64) between the density matrix and the resolvent, we expect that this solution predicts the depth of the resonance band in the complex plane of $k$.
Furthermore, the transport equation (72) is equivalent to the linear Boltzmann equation for the Wigner function of the quantum particle [25, 33, 34], as shown in Appendix G. This reveals an interesting connection with decoherence theory of a quantum particle in a gas of moving atoms [74, 75, 76, 77, 78]. In the Foldy-Lax model, where the scatterers are fixed and the disorder is quenched, the randomness of the scatterer positions leads to decoherence, which is characterized by a finite coherence length [79, 80, 77, 78]
$\Lambda(\boldsymbol{\mathrm{r}}_{0})^{2}=\frac{\int_{\mathbb{R}^{d}}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0}\right\|^{2}\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}_{0})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}}{\int_{\mathbb{R}^{d}}\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}_{0})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}}\>.$ (74)
The length (74) is nothing but the standard deviation of the distribution $\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\right|^{2}$ around a reference point, say $\boldsymbol{\mathrm{r}}^{\prime}$. This is a measure of the characteristic length beyond which the statistical correlations are lost. Indeed, the average density matrix is also the autocorrelation function of the wavefunction, i.e., $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\rangle=\langle\psi(\boldsymbol{\mathrm{r}}){\psi}^{*}(\boldsymbol{\mathrm{r}}^{\prime})\rangle$. Beyond the coherence length, the wave formalism becomes less meaningful in the sense that two wave packets located far enough from each other are statistically independent and experience different collision histories without possible interference between them.
Let us determine the coherence length (74) in the special case of an infinite disordered medium and at equilibrium, that is after a time much longer than the mean intercollisional time. It is shown in Appendix H that, in this regime, the density matrix obeying Eq. (68) asymptotically behaves as
$\begin{split}\rho_{\infty}(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\xrightarrow{r\rightarrow\infty}&\frac{1}{2k_{0}}\left(\frac{k_{0}}{2\pi r}\right)^{\frac{d-1}{2}}\\ &\times\sin\left(k_{0}r-\tfrac{\pi(d-3)}{4}\right)\mathop{}\!\mathrm{e}^{-\frac{n\sigma}{2}r}\>,\end{split}$ (75)
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$. Substituting Eq. (75) into Eq. (74) provides the sought coherence length
$\Lambda(\boldsymbol{\mathrm{r}}_{0})=\sqrt{2}\,\ell_{\rm s}\>.$ (76)
This result shows that the coherence length is of the order of the mean free path. In this model, the mean free path is thus not only the mean distance between two successive collisions, but also the coherence length. Due to the analogy with decoherence theory [74, 75, 76, 81, 82, 77, 78] mentioned before, the result (76) confirms the interpretation of the exponential evanescence of the average Green function at the scale of the mean free path highlighted in Eq. (46) as being related to decoherence.
IV.3 Diffusion approximation
The issue with the transport equation (72) is that solving it analytically is rarely feasible, even in simple geometries such as the spherical geometry considered in this paper. In addition, this equation is quite difficult to solve numerically because it generally requires too large discretization lattices to be practical. However, it is possible to considerably simplify the problem in the diffusion regime, that is when the mean free path $\ell_{\rm s}$ is much smaller than the radius $R$ of the spherical medium.
In this section, we compare the average density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ predicted in the diffusion approximation to the actual density of the wavefunction for both a single realization of the disorder and on average. In addition, we show that the diffusion approximation leads to an accurate estimate of the depth of the resonance band in Figs. 3(a) and 3(b) with respect to the real $k$ axis for $d\geq 2$.
As shown in Appendix I, Eq. (72) reduces in the diffusion regime and far away from the boundaries to
$\gamma\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=(\gamma+n\sigma)K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+\frac{n\sigma}{d(\gamma+n\sigma)^{2}}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle\>.$ (77)
Note that Eq. (77) is not exactly the standard diffusion equation because the Laplace variable $\gamma$, which can be interpreted as a time derivative ($\gamma\leftrightarrow\partial_{vt}$), is not only in the left-hand side, but also in the right-hand side. However, since we are only interested in time scales much longer than the intercollisional time ($t\gg(n\sigma v)^{-1}$), the variable $\gamma$ is much smaller than $n\sigma$ and can be neglected in the right-hand side of Eq. (77). Then, using the property $(\gamma+n\sigma)K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\simeq\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})$ in the diffusive regime, Eq. (77) becomes
$\gamma\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})+\frac{1}{dn\sigma}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle\>.$ (78)
To let appear the diffusion equation out of Eq. (78), we have to perform the inverse Laplace transform in time
$\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\frac{1}{2\pi\mathrm{i}}\int_{-\mathrm{i}\infty}^{\mathrm{i}\infty}\mathop{}\!\mathrm{e}^{\gamma vt}\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle\mathop{}\!\mathrm{d}(\gamma v)\>,$ (79)
recalling that $s=\gamma v$. This is why, in Eq. (78), we have stressed the dependence of $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ on $\gamma$. We ultimately obtain the diffusion equation
$\frac{\partial}{\partial t}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle-D\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})\delta(t)\>,$ (80)
where the parameter $D$ is the diffusivity, or diffusion constant, characterizing the diffusive transport of the particle in the medium. It reads
$D=\frac{v}{dn\sigma}=\frac{v\ell_{\rm s}}{d}\>.$ (81)
Note that the right-hand side of Eq. (80) contains a source term at $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{r}}_{0}$ at time $t=0$. This is consistent with the assumptions made in Sec. IV.1 about the initial wave packet $\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|$. The diffusion equation (80) can be solved in an infinite medium. Setting $\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$, the result is [25]
$\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\frac{1}{(4\pi Dt)^{\frac{d}{2}}}\exp\left(-\frac{r^{2}}{4Dt}\right)\Theta(t)\>,$ (82)
where $\Theta(t)$ denotes the Heaviside step function. Expression (82) is a normalized Gaussian distribution. The variance of this distribution is given by
$\left\langle\boldsymbol{\mathrm{r}}(t)^{2}\right\rangle=\int_{\mathbb{R}^{d}}\left\|\boldsymbol{\mathrm{r}}\right\|^{2}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}=2dDt=2\ell_{\rm s}vt\>,$ (83)
and thus linearly increases in time.
The diffusion equation must be supplemented by the condition on the spherical boundary ($\forall\boldsymbol{\mathrm{r}}\in\mathcal{S}_{d}(R)$),
$\left(\gamma+\frac{n\sigma}{2}\right)\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\frac{n\sigma}{\gamma+n\sigma}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>,$ (84)
which is derived at the end of Appendix I. When considering times much longer than the intercollisional time, the variable $\gamma$ in Eq. (84) can be neglected in front of $n\sigma$. In this way, we obtain a time-independent version of Eq. (84), or rather a $\gamma$-independent version:
$\frac{n\sigma}{2}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>.$ (85)
This is a Robin boundary condition on the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ for all $\boldsymbol{\mathrm{r}}\in\mathcal{S}_{d}(R)$. Since the coefficients of the two terms in Eq. (85) are positive, this condition expresses the fact that the density has an inward-facing gradient. In the far diffusive regime ($n\sigma R\rightarrow\infty$), this condition reduces to a simple Dirichlet boundary condition: $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0$. Although this regime is valid for a very large medium, we do not consider this extreme case in this paper.
IV.4 Spatial profile of the wavefunction
In this section, we compare the predictions of the diffusion equation regarding the average density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ to the numerical results obtained in the Foldy-Lax model with the program MSModel [56]. In particular, we want to know if the disordered-average density, $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$, is representative of the density in a single realization of the disorder. In other words, we wonder if the system has some self-averaging property.
Furthermore, we want to compare the prediction of the diffusion equation (78) to the numerical value of $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$ in the spherical-shaped point field. For this purpose, we assume that the wave packet is created at the center of the spherical medium ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$). In the stationary regime ($\gamma\rightarrow 0$), the Laplace-transformed diffusion equation (78) reads
$\nabla^{2}_{\boldsymbol{\mathrm{r}}}\rho_{\rm da}(\boldsymbol{\mathrm{r}})=-A\,\delta(\boldsymbol{\mathrm{r}})\>,$ (86)
where the subscript stands for diffusion approximation, and the amplitude factor is $A=dn\sigma$. However, the normalization convention in our numerical simulations is slightly different from what Eq. (78) prescribes. In these simulations, we assume that the source term in Eq. (72) is $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$ and not $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$. The correction is simply given by the prefactor in Eq. (70). Thus, we should have
$A=dn\sigma\frac{\pi\mu(k_{0})}{k_{0}}\>.$ (87)
The general solution of the Poisson equation (86) reads
$\rho_{\rm da}(r)=\begin{cases}-\frac{A}{S_{d}}\frac{r^{2-d}-R_{\rm eff}^{2-d}}{2-d}&\text{for}~{}d\neq 2\>,\\[10.0pt] -\frac{A}{2\pi}\ln\left(\frac{r}{R_{\rm eff}}\right)&\text{for}~{}d=2\>,\end{cases}$ (88)
where the arbitrary parameter $R_{\rm eff}$ denotes the effective radius at which the density is supposed to vanish as if the medium was extended beyond $R$. Since the density must be positive everywhere, the effective radius is necessarily larger or equal to the actual radius $R$ of the medium ($R_{\rm eff}\geq R$). The effective radius has to be adjusted to satisfy the boundary condition (85), which can also be rewritten as
$\frac{\rho_{\rm da}(R)}{\partial_{r}\rho_{\rm da}(R)}=-\frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\>.$ (89)
In order to satisfy Eq. (89), the effective radius $R_{\rm eff}$ in Eq. (88) must read
$R_{\rm eff}=\begin{cases}R\left(1+(2-d)\frac{2V_{d-1}}{S_{d}}\frac{\ell_{\rm s}}{R}\right)^{\frac{1}{2-d}}&\text{for}~{}d\neq 2\>,\\[10.0pt] R\exp\left(\frac{2V_{d-1}}{S_{d}}\frac{\ell_{\rm s}}{R}\right)&\text{for}~{}d=2\>,\end{cases}$ (90)
Note that, if the mean free path is small enough ($\ell_{\rm s}\ll R$), the expressions (90) can be effectively expanded at first order. We find
$R_{\rm eff}\simeq R+\frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\>.$ (91)
Since the dimensionless factor $2V_{d-1}/S_{d}$ is of the order of one, the result (91) means that the effective radius is about the actual radius plus the mean free path. This is a known result of transport theory, which is, for instance, encountered in the neutronics of nuclear reactors [34].
The square modulus of the wavefunction numerically computed with the program MSModel [56] in a spherical-shaped random point field for $N=10^{3}$ is shown in Fig. 4 and the cross-sectional view long the $x$ axis in Fig. 5. The similar cross-sectional view of Fig. 1(a) is shown in Fig. 6.
Figure 4: (a) Square modulus of the wavefunction computed with the program MSModel [56] in a 2D disk-shaped random point field of $N=10^{3}$ scatterers with $k=10\,\varsigma^{-1}$. The point scattering model is Eq. (15) and the mean free path is $\ell_{\rm s}=2.5\,\varsigma$. The point source is located at the center ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$). The circle depicts the medium boundary. (b) Same as panel (a) but averaged over $2^{6}$ random configurations.
Figure 5: (a) Cross-sectional view of Fig. 4(a) along the positive $x$ semiaxis. The vertical gray line indicates the radius of the medium. The dashed curve is the diffusion approximation (88) using Eqs. (87) and (90). (b) Cross-sectional view of Fig. 4(b) averaged over $2^{14}$ configurations. The filled region is the numerically computed interquartile range of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$. The dotted curve is the square modulus of the effective Green function (40).
Figure 6: Cross-sectional view of Fig. 1(a) along the positive $x$ semiaxis. The vertical gray line indicates the radius of the medium. The dashed curve is the diffusion approximation (88) using Eqs. (87) and (90).
We note that the filament-like structures observed in Fig. 1(a) and Fig. 4(a) have a scale of the order of the mean free path $\ell_{\rm s}$ introduced by Eq. (44) in relation to the free flights of the particle between its collisions on the scatterers.
Figure 4(b) shows the wavefunction density averaged over several random configurations of the scatterers. We observe that the detailed structures of the wavefunction seen in Fig. 4(a) and, in particular, the random meandering nodal lines have disappeared after averaging.
In both Figs. 5 and 6, the dashed curve is the theoretical average density (88) predicted by the diffusion approximation. The matching with the numerical average for $r\leq R$ in Fig. 5(b) is remarkable, despite the apparent roughness of the diffusion approximation. Indeed, the scale parameter $R/\ell_{\rm s}\simeq 7$ is not so large compared to one. Figures 5(a) and 6 seem to confirm that the wavefunction in a single configuration globally follows the theoretical decrease of $\rho_{\rm da}(r)$. This suggests that the aspect of the wave density in an infinite disordered medium will be closer to Fig. 4(b) than to Fig. 4(a), because the medium will ultimately look uniform at such large scales. This would support the validity of the disorder averaging for a sufficiently large medium ($R\gg\ell_{\rm s}$).
In the neighborhood of the source point at $\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$, the numerical average does not match the diffusion approximation in Fig. 5(b). The reason is that the diffusion approximation is not supposed to be accurate on scales shorter than the mean free path. Indeed, the amplitude of the singularity predicted by the diffusion approximation is not consistent with that of the Bethe-Salpeter equation (72). According to Eq. (68), the intensity near the source point $\boldsymbol{\mathrm{r}}_{0}$ is approximately given by the effective Green function $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$. This is confirmed by the matching between the solid and the dotted curves near $r=0$ in Fig. 5(b).
Besides that, one notices a mismatch at the edge of the medium for $r\geq R$. This is due to the fact that the density (88) is only valid within the point field. In the outer region ($r\geq R$), the density behaves as the power law
$\rho(r)\propto\frac{1}{r^{d-1}}\>.$ (92)
This is given by the far-field behavior of Eq. (72) for $r\gg R$.
IV.5 Depth of the resonance band
The integral equation (72) also provides a prediction of the position of the band of complex resonances seen in Figs. 3(a) and 3(b) but also in our previous paper [47]. The resonant states can be obtained as the eigenvectors of the Schrödinger equation without any source term, just like eigenstates but at complex energies. In the same way, in order to study the resonance band from Eq. (72), we need to set the source term $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ to zero and interpret the resulting equation
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=\int_{\mathbb{R}^{d}}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (93)
as an eigenvalue problem, where the free variable $\gamma$ plays the role of the sought eigenvalue and the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ is the associated eigenfunction. The eigensystem (93) is more easily solved in the diffusion approximation. Using Eqs. (78) and (85), the integral equation (93) can be cast into
$\begin{cases}\frac{\ell_{\rm s}}{d}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle-\gamma\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}r\leq R\>,\\[10.0pt] \frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0&\text{for}~{}r=R\>.\end{cases}$ (94)
The first line of Eq. (94) is the Laplace-transformed diffusion equation, and the second line is the boundary condition at $r=R$. Indeed, we assume that the medium is spherical. The first line of Eq. (94) can be written in the more familiar form
$(\nabla^{2}_{\boldsymbol{\mathrm{r}}}+\beta^{2})\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>,$ (95)
where $\beta$ is the Fourier variable for space, analogous to a wavenumber. The solution of Eq. (95) reads
$\left\langle\rho(r)\right\rangle=j^{(d)}_{0}(\beta r)\>,$ (96)
where $j^{(d)}_{\ell}(z)$ is the generalized spherical Bessel function defined in Eq. (116). It mainly depends on the Bessel function $J_{\frac{d-2}{2}}(z)$. Using Eq. (96), the differential problem (94) becomes the following nonlinear system for the pair $(\gamma,\beta)$
$\begin{cases}\frac{\ell_{\rm s}}{d}\beta^{2}+\gamma=0\>,\\[10.0pt] \frac{2V_{d-1}}{S_{d}}\beta\ell_{\rm s}{j^{(d)}_{0}}^{\prime}(\beta R)+j^{(d)}_{0}(\beta R)=0\>.\end{cases}$ (97)
The system of equations (97) is graphically solved in Fig. 7(a).
Figure 7: (a) Graphical solution of the diffusion eigenmodes from Eq. (97) in the dual space $(\gamma,\beta)$ for a 2D disk-shaped random point field with $N=500$ scatterers. The parameters are $R=12.6\,\varsigma$ and $\ell_{\rm s}=1.5\,\varsigma$. The black dots are the first four solutions: $n\in\{1,2,3,4\}$. (b) First diffusion eigenmodes $n\in\{1,2,3\}$ in the position space from Eq. (96).
The first equation corresponds to the parabola, and the second equation to the horizontal straight lines. The intersections between both are the eigenmodes of Eq. (94). The corresponding eigenvalues, denoted as $\beta_{n}$ and $\gamma_{n}$, fully characterize the eigenmodes in space and time, respectively. The profiles of the first diffusion eigenmodes in space are shown in Fig. 7(b) based on Eq. (96).
Furthermore, the spectral decomposition of the diffusion problem (94) allows us to write the general solution for the density and its time evolution. Through inverse Laplace transform, one has the expansion
$\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\sum_{n=1}^{\infty}c_{n}\mathop{}\!\mathrm{e}^{\gamma_{n}vt}j^{(d)}_{0}(\beta_{n}r)\>,$ (98)
where $v$ is the group velocity and $c_{n}$ are arbitrary constants depending on the initial condition at $t=0$. Since the density must be positive, the constant $c_{1}$, associated with the fundamental mode, should be the largest one. Note that the values of $\gamma_{n}$ are negative. Therefore, Eq. (98) describes the decay of the density in time due to the escape of the particle from the medium. In the long time limit ($t\rightarrow\infty$), the fundamental mode dominates the other ones, since $\gamma_{1}$ has the smallest value. Of course, the superiority of the fundamental mode is required by the constraint on the positivity of the density.
Although the boundary condition in Eq. (97) cannot be solved for $\beta$ in closed form, it is possible to continue the calculation further if the mean free path is sufficiently small ($\ell_{\rm s}\ll R$). Indeed, for a small enough $\ell_{\rm s}$, one may recognize the first order expansion of
$j^{(d)}_{0}\!\left(\beta R+\tfrac{2V_{d-1}}{S_{d}}\beta\ell_{\rm s}\right)=0\>,$ (99)
in the second equation of Eq. (97). Using the approximate effective radius (91), Eq. (99) can be simplified to
$j^{(d)}_{0}(\beta R_{\rm eff})=0\>.$ (100)
Condition (100) obviously expresses the geometrical interpretation of $R_{\rm eff}$ as the point where the extrapolated density vanishes. This interpretation was already discussed near Eq. (88). The roots of Eq. (100) are given by the zeros of $j^{(d)}_{0}(z)$, or, more explicitly, by the zeros of $J_{\frac{d-2}{2}}(z)$. Thus, we obtain the full spectrum of diffusion eigenmodes
$\beta_{n}=\frac{j_{\frac{d-2}{2},n}}{R_{\rm eff}}\>,$ (101)
where $j_{\nu,n}~{}\forall n\in\{1,2,3,\ldots\}$ denotes the $n$-th zero of the Bessel function $J_{\nu}(z)$. The first zeros of interest are given in Tab. 2 for the lowest dimensions.
$d$ 1 2 3 4
$n=1$ $\tfrac{\pi}{2}$ $2.40483\ldots$ $\pi$ $3.83171\ldots$
$n=2$ $\tfrac{3\pi}{2}$ $5.52008\ldots$ $2\pi$ $7.01559\ldots$
$n=3$ $\tfrac{5\pi}{2}$ $8.65373\ldots$ $3\pi$ $10.1735\ldots$
$n=4$ $\tfrac{7\pi}{2}$ $11.7915\ldots$ $4\pi$ $13.3237\ldots$
$n=5$ $\tfrac{9\pi}{2}$ $14.9309\ldots$ $5\pi$ $16.4706\ldots$
Table 2: First zeros of the Bessel function $J_{\frac{d-2}{2}}(z)$, denoted as $j_{\frac{d-2}{2},n}$ in the text [55]. No closed form exists for these zeros, except for $d=1$ and $d=3$.
According to the first line of Eq. (97), the values of $\gamma$ associated to Eq. (101) read
$\gamma_{n}=-\frac{\ell_{\rm s}}{d}\left(\frac{j_{\frac{d-2}{2},n}}{R_{\rm eff}}\right)^{2}\>.$ (102)
The result (102) turns out to be an accurate approximation of the sought eigenvalue of Eq. (93). The quantity (102) is also related to the escape rate of the particle from the medium by $\Gamma_{n}=\left|\gamma_{n}\right|v$.
Finally, using the relationship $\gamma=2{k}_{\rm i}$ introduced in Eq. (45), the diffusion rate of Eq. (102) translates into the following estimate
$k_{{\rm idiff},n}=-\frac{\ell_{\rm s}}{2d}\left(\frac{j_{\frac{d-2}{2},n}}{R_{\rm eff}}\right)^{2}\>,$ (103)
for the position of the resonance band. The result (103) is considerably more accurate than Eq. (68) of Ref. [47] since it accounts for the Robin boundary condition (97) which is more precise than the Dirichlet condition $j^{(d)}_{\ell}(\beta R)=0$. Equation (103) is also more general than Eq. (68) because it predicts other diffusion eigenmodes than the fundamental one.
The comparison between the prediction (103) and the numerical results is shown in Fig. 8 for the 2D and 3D cases.
Figure 8: (a) Cross-sectional view of the resonance density for a 2D disk-shaped random point field with $N=500$ along the vertical axis ${k}_{\rm r}=6\,\varsigma^{-1}$ and averaged over $2^{10}$ configurations. The parameters are $R=12.6\,\varsigma$ and $\ell_{\rm s}=1.5\,\varsigma$. The vertical dotted lines indicate the diffusion eigenmodes given by Eq. (97). (b) Same as panel (a) but in the 3D case with $R=4.92\,\varsigma$ and $\ell_{\rm s}=2.865\,\varsigma$.
The resonance density (solid curve) is computed with the program MSModel [56] exploiting the resonance potential method of our previous paper [47]. Note that the final increase of the density in Fig. 8(a) near ${k}_{\rm i}=1\,\varsigma^{-1}$ is due to numerical round-off errors in computing the determinant of $\mathsf{M}(k)$ and is thus not physical. This phenomenon was discussed in our previous paper [47].
The adequacy between the fundamental diffusion mode ($n=1$) and the maximum of the resonance band in Fig. 8 is remarkable. This is especially true in 3D, because the scale parameter $n\sigma R=1.72$ is not much larger than 11 11 compared to what the diffusion approximation discussed in Sec. IV.3 could have required. In fact, the quality of the approximation (103) greatly depends on the expression of $R_{\rm eff}$ and on the boundary condition. Indeed, if the magnitudes of $R$ and $\ell_{\rm s}$ are the same, there is a factor of two between $R$ and $R_{\rm eff}$, and thus a factor of four in $k_{{\rm idiff},n}$, which is significant. This shows that the accuracy of the prediction (103) is not obvious.
Noticeably, the other diffusion eigenmodes do not lead to any perceptible structure in Fig. 8. We could have thought that Eq. (103) predicts the existence of additional bands deeper in the complex plane. However, we could not find any match, not even in much larger point fields, up to $N=5000$. Only the fundamental mode ($n=1$) seems to play a significant role. Furthermore, regarding the 1D case, these higher-order modes do no coincide with the faint bands in Fig. 6(a) of our previous paper [47]. All these observations lead us to conjecture that these higher-order diffusion eigenmodes do not exist in the quantum case.
A likely reason of this absence is that the transport equation (72), on which Eq. (103) is based, only describes the incoherent (phase-independent) contribution to wave transport. Therefore, this equation is unable to predict the full resonance distribution in the complex $k$ plane, because the latter would be intrinsically determined by coherence and constructive interferences in the same way as the energy eigenstates in quantum mechanics. If this hypothesis is confirmed, the effect of coherent contribution would be to smooth out the comb of higher-order diffusion eigenmodes in Fig. 8.
The issue of the resonance distribution could be overcome with advanced methods capable of describing the coherent contribution to wave transport such as the nonlinear sigma model [83]. This field-theoretical method is based on the average of some generating function over the realizations of a random potential. Upon averaging, the linear wave equation in the disordered medium is then cast into a disorder-free nonlinear wave equation which can be solved semiclassically. This method has proven particularly effective in determining the full statistics of certain observables [83, 84, 43].
V Conclusion and perspectives
In this paper, we studied the Foldy-Lax model to describe the propagation of a quantum particle in a disordered medium such as a gaseous detector. In this model, the atoms composing the medium are represented by point scatterers pinned in space at random positions. We studied the properties of this system under the perspective of the resonance poles in the complex plane of the wavenumber $k$. We developed equations for disorder-averaged quantities, in particular the wavefunction and the distribution of complex resonances. These equations were compared to numerical simulations performed with the program MSModel [56].
More specifically, we showed in Sec. III that the peaks in the resonance density at low energy can be predicted by an effective-medium equation for the disorder-averaged wavefunction. This means that, when the particle wavelength is larger than the mean interscatterer distance, the wave only probes the geometric shape of the medium and not the details due to the disorder.
When the wavelength is smaller than the interscatterer distance, the Bethe-Salpeter equation is required to describe the multiple scattering in the medium. In Sec. IV.4, we checked that the average square modulus of the wavefunction in the bulk predicted by the Bethe-Salpeter equation is consistent with numerical simulations. The square modulus of the wavefunction corresponding to a single configuration of scatterers was also considered (see in particular Fig. 1 for $N=10^{4}$ scatterers) in order to determine the extent of fluctuations relative to the average. These simulations revealed the presence of filaments in the square modulus of the wavefunction which we interpreted as a phenomenon similar to quantum scars [61, *Heller2001]. The detailed study of these structures, including their possible links with the multiple-scattering resonances, is deferred to future works.
Finally, in Sec. IV.5, we solved the Bethe-Salpeter equation in the diffusion regime in order to estimate the resonance band. We showed that the fundamental diffusion eigenmode precisely corresponds to the imaginary part of the resonance band, thus confirming that this band is related to the escape time of the particle in the medium. However, the higher-order diffusion modes given by the diffusion equation do not lead to visible structures in the resonance density. A likely reason is that the Bethe-Salpeter equation ignore coherent effects, although they are crucial in determining energy eigenstates. In the future, we plan to use advanced methods accounting for coherent contributions to wave transport, such as the nonlinear sigma model [83, 84, 43], in order to finely characterize the distribution of resonance widths in strongly open systems.
Acknowledgements.
The present results were obtained during D.G.’s doctoral thesis defended in June 2022. The authors are grateful to Pierre Gaspard for useful discussions and for reviewing this manuscript. The authors also thank Arthur Goetschy and Romain Pierrat for their support in this research. D.G. is grateful to Arthur Goetschy for bringing Refs. [83, 84] to his attention. This work was supported by the Belgian National Fund for Scientific Research (F.R.S.-FNRS) as part of the “Research Fellow” (ASP - Aspirant) fellowship program. This work was also supported by the F.R.S.-FNRS as part of the Institut Interuniversitaire des Sciences Nucléaires (IISN) under Grant No. 4.45.10.08. It also received funding from the French National Research Agency (ANR) as part of the “MARS LIGHT” project.
Appendix A Derivation of the disorder-averaged Green function
We present here a derivation of the effective-medium equation that we consider for the disorder-averaged Green function using the diagrammatic method. The matrix $\mathsf{M}(k)^{-1}$ in Eq. (28) can be expanded in Dyson series as
$\hat{Q}(k)=\sum_{p=0}^{\infty}\sum_{i_{1},\ldots,i_{p}}\hat{G}(k)\hat{F}_{i_{p}}\hat{G}(k)\cdots\hat{G}(k)\hat{F}_{i_{2}}\hat{G}(k)\hat{F}_{i_{1}}\hat{G}(k)\>,$ (104)
where $\hat{F}_{i}=F(k)\left|\boldsymbol{\mathrm{x}}_{i}\right\rangle\left\langle\boldsymbol{\mathrm{x}}_{i}\right|$ denotes the projector onto the $i$-th scatterer weighted by the single-scatterer amplitude. The term $p=0$ in Eq. (104) corresponds to the free propagation: $\hat{G}(k)$. Note that two successive indices, for instance $i_{1}$ and $i_{2}$, cannot be equal to each other, since the Green matrix $\mathsf{G}^{+}(k)$ in Eq. (21) is zero along the diagonal. However, terms such as $i_{1}=i_{3}$ or $i_{1}=i_{4}$ are allowed, for instance. An elegant way of visualizing and classifying the different kinds of terms involved in Eq. (104) is to resort to diagrammatic representations. We consider for this purpose the diagrammatic notations given in Fig. 9(a) and inspired by Refs. [27, 52, *Doicu2019a2, *Doicu2019a3].
Figure 9: (a) Diagrams for the free Green function and the scattering vertex. (b) Expansion of the full resolvent $\hat{Q}(k)$ from Eq. (104). The list of diagrams is not exhaustive. The sums implicitly assume that the indices $i_{1},i_{2},\ldots$ are different from each other. Pairs of identical scatterers are highlighted by a black dotted link.
With these notations, some lowest order terms of the series (104) are shown in Fig. 9(b). In our convention, the diagrams should be read from right to left in the chronological order of the scattering events. Of course, this is not essential because of the time reversal symmetry of the Schrödinger equation. The black dotted links in Fig. 9(b) stress the fact that the scatterer encountered by the particle is the same. The existence of links between adjacent scatterers is forbidden by the zero on the diagonal of $\mathsf{G}(k)$. This is because the collision of the particle on a single scatterer is assumed to be exactly described by the scattering amplitude $F(k)$ and does not need any additional correction.
The full zoo of diagrams in Fig. 9(b) is very difficult to treat in general, especially the great herd of linked diagrams. In order to deal with the series (104), we have to neglect those linked diagrams. This approximation is sometimes known as the Twersky approximation [63, 27, 52, *Doicu2019a2, *Doicu2019a3], although it was already present in Foldy’s paper [28]. A similar approximation in the context of continuous disorder is due to Bourret [68, 19, 20, 21]. The Twersky approximation is based on the idea that, for a sufficiently large set of scatterers ($N\rightarrow\infty$), the chance that the particle collides two times the same scatterer is negligible. The only diagrams which are taken into account consider that the particle takes self-avoiding paths. In this regard, the sums over $i_{1},i_{2},i_{3},\ldots$ in Fig. 9(b) implicitly assume that the indices are all different from each other.
One may verify that the number of neglected diagrams is relatively small for $N\rightarrow\infty$. Indeed, at a given order $p$, the total count of diagrams in Fig. 9(b) is $\mathcal{C}_{\rm tot}=N(N-1)^{p-1}$, whereas the count reduces to $\mathcal{C}_{\rm sa}=N!/(N-p)!$ if we limit ourselves to self-avoiding paths under the Twersky approximation [27, 52, *Doicu2019a2, *Doicu2019a3]. The relative difference between these counts behaves as
$\frac{\mathcal{C}_{\rm tot}-\mathcal{C}_{\rm sa}}{\mathcal{C}_{\rm sa}}=\frac{(p-1)(p-2)}{2N}+\operatorname{\mathcal{O}}(p^{4}/N^{2})\>,$ (105)
which obviously tends to zero for $N\rightarrow\infty$. Equation (105) confirms that the diagrams for which the particle visits the same scatterer several times are less numerous than the self-avoiding paths ($\mathcal{C}_{\rm tot}-\mathcal{C}_{\rm sa}\ll\mathcal{C}_{\rm sa}$). This supports the validity of the Twersky approximation under the large $N$ assumption [27, 52, *Doicu2019a2, *Doicu2019a3].
The Twersky approximation considerably simplifies the calculation of the average Green operator $\langle\hat{Q}(k)\rangle$. Indeed, since the positions of the scatterers are supposed to be statistically independent in our model, the configurational averages can be factored as follows
$\begin{split}&\left\langle\hat{G}(k)\sum_{i_{1}}\hat{F}_{i_{1}}\hat{G}(k)\cdots\hat{G}(k)\sum_{i_{p}}\hat{F}_{i_{p}}\hat{G}(k)\right\rangle\\ &=\hat{G}(k)\left\langle\sum_{i_{1}}\hat{F}_{i_{1}}\right\rangle\hat{G}(k)\cdots\hat{G}(k)\left\langle\sum_{i_{p}}\hat{F}_{i_{p}}\right\rangle\hat{G}(k)\\ &=\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\cdots\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\>,\end{split}$ (106)
where one has used several times the result
$\left\langle\sum_{i=1}^{N}\hat{F}_{i}(k)\right\rangle=\left\langle\sum_{i=1}^{N}F(k)\delta(\hat{\boldsymbol{\mathrm{r}}}-\boldsymbol{\mathrm{x}}_{i})\right\rangle=n(\hat{\boldsymbol{\mathrm{r}}})F(k)\>,$ (107)
coming from Eq. (5). Note that $n(\hat{\boldsymbol{\mathrm{r}}})$ does not commute with $\hat{G}(k)$ in general, because the point field has a finite size and the density thus depends on the position $\hat{\boldsymbol{\mathrm{r}}}$. Substituting Eq. (106) into Eq. (104) gives us the average Green operator
$\begin{split}\langle\hat{Q}(k)\rangle&=\hat{G}(k)\\ &+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\\ &+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\hat{G}(k)\\ &+\cdots\>.\end{split}$ (108)
Again, this series can be more easily visualized using diagrams. We introduce to this end the additional notations given in Fig. 10(a) and inspired by Ref. [53].
Figure 10: (a) Diagrams for the effective Green function and the effective vertex. (b) Resummation of the diagrams of the average resolvent, $\langle\hat{Q}(k)\rangle$, under the Twersky approximation. This step is equivalent to Eqs. (108) and (109).
With these notations, the series (108) translates to the first row of Fig. 10(b). In particular, it becomes more obvious that the series can be cast onto the recursive form
$\langle\hat{Q}(k)\rangle=\hat{G}(k)+\hat{G}(k)n(\hat{\boldsymbol{\mathrm{r}}})F(k)\langle\hat{Q}(k)\rangle\>.$ (109)
With a little bit of algebra, Eq. (109) can also be rewritten as
$\langle\hat{Q}(k)\rangle=\frac{1}{\hat{G}(k)^{-1}-n(\hat{\boldsymbol{\mathrm{r}}})F(k)}=\frac{1}{k^{2}-\hat{\boldsymbol{\mathrm{p}}}^{2}-n(\hat{\boldsymbol{\mathrm{r}}})F(k)}\>.$ (110)
Finally, Eq. (110) directly provides the effective-medium equation (39).
Appendix B Scattering by a sphere and resonances
In this appendix, we derive the formula (49) for the complex resonance poles corresponding to the scattering by a spherical effective medium. The wave equation is Eq. (38) with the spherical-shaped density
$n(\boldsymbol{\mathrm{r}})=\begin{cases}n&\text{if}~{}\left\|\boldsymbol{\mathrm{r}}\right\|\leq R\>,\\ 0&\text{otherwise}\>,\end{cases}$ (111)
where $R$ is the radius of the disordered medium. The effective wavenumber in the region $\left\|\boldsymbol{\mathrm{r}}\right\|\leq R$ is thus given by Eq. (41). In order to solve Eq. (38) given the spherical symmetry of Eq. (111), we will use the partial wave method developed in Sec. 1.2 of Ref. [67]. In this method, the average wavefunction can be expanded in partial waves as
$\left\langle\psi(\boldsymbol{\mathrm{r}})\right\rangle=\sum_{\ell=0}^{\infty}\left\langle\psi_{\ell}(r)\right\rangle P^{(d)}_{\ell}(\cos\theta)\>,$ (112)
where $P^{(d)}_{\ell}(\cos\theta)$ denotes the Gegenbauer polynomials generalizing the Legendre polynomials to arbitrary dimension $d$ and defined by
$P^{(d)}_{\ell}(x)=\frac{\Gamma(\ell+d-2)}{\ell!\,\Gamma(d-1)}{}_{2}F_{1}\!\left(\begin{subarray}{c}-\ell,\,\ell+d-2\\ \frac{d-1}{2}\end{subarray};\frac{1-x}{2}\right)\>,$ (113)
where ${}_{2}F_{1}\!\left(\begin{subarray}{c}a,\,b\\ c\end{subarray};z\right)$ is the Gauss hypergeometric function [55]. In Eq. (112), $\cos\theta=\boldsymbol{\mathrm{\Omega}}\cdot\boldsymbol{\mathrm{\Omega}}_{0}$ is the cosine of the angle of $\boldsymbol{\mathrm{\Omega}}=\boldsymbol{\mathrm{r}}/r$ with respect to the direction $\boldsymbol{\mathrm{\Omega}}_{0}$ of the incident plane wave. The radial components of the average wavefunction in Eq. (112) satisfy the radial Schrödinger equation
$\begin{split}\bigg{[}\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}r^{2}}+\frac{d-1}{r}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}r}+k^{2}&-\frac{\ell(\ell+d-2)}{r^{2}}\\ &-n(r)F(k)\bigg{]}\left\langle\psi_{\ell}(r)\right\rangle=0\>,\end{split}$ (114)
where the density is given by Eq. (6) for a spherical-shaped medium of radius $R$. According to this expression of the density, we have to split the wavefunction $\langle\psi_{\ell}(r)\rangle$ in two regions $r\leq R$ and $r>R$. In this regard, we consider the following ansatz
$\left\langle\psi_{\ell}(r)\right\rangle=\begin{cases}j^{(d)}_{\ell}(\kappa r)&\text{for}~{}r\leq R\>,\\ S_{\ell}(k)h^{+(d)}_{\ell}(kr)+h^{-(d)}_{\ell}(kr)&\text{for}~{}r>R\>,\end{cases}$ (115)
where the generalized spherical Bessel functions are defined by
$h^{\pm(d)}_{\ell}(z)$ $=\Gamma(\tfrac{d}{2})\left(\frac{2}{z}\right)^{\frac{d-2}{2}}H^{\pm}_{\ell+\frac{d-2}{2}}(z)\>,$ (116)
$j^{(d)}_{\ell}(z)$ $=\frac{h^{+(d)}_{\ell}(z)+h^{-(d)}_{\ell}(z)}{2}\>,$
where $H^{\pm}_{\nu}(z)$ are the Hankel functions [55]. The function $j^{(d)}_{\ell}(z)$ is finite and regular at $z=0$, hence its presence in the inner region of the ansatz (115). Imposing the continuity condition on the wavefunction and its derivative at the boundary $r=R$, we quickly get the following expression for the scattering matrix element
$S_{\ell}(k)=-\frac{\operatorname{W}[h^{-(d)}_{\ell}(kr),j^{(d)}_{\ell}(\kappa r)]_{r=R}}{\operatorname{W}[h^{+(d)}_{\ell}(kr),j^{(d)}_{\ell}(\kappa r)]_{r=R}}\>,$ (117)
where $\operatorname{W}[f(r),g(r)]=f(r)\partial_{r}g(r)-g(r)\partial_{r}f(r)$ denotes the Wronskian with respect to $r$.
We look for the resonance poles in the scattering matrix element (117) coming from the cancellation of the denominator
$\operatorname{W}[h^{+(d)}_{\ell}(kr),j^{(d)}_{\ell}(\kappa r)]_{r=R}=0\>.$ (118)
Finally, expressing $j^{(d)}_{\ell}(z)$ and $h^{+(d)}_{\ell}(z)$ in terms of the standard Bessel and Hankel functions, $J_{\nu}(z)$ and $H^{+}_{\nu}(z)$, we retrieve the resonance equation (49).
Appendix C Zeros of the Hankel function
Here, we study the resonance spectrum corresponding to the hard-sphere scattering, which is given by the complex zeros of the Hankel function $H^{+}_{\nu}(z)$, i.e., Eq. (51). As discussed in Sec. III.3, these zeros turn out to be very close to the resonance poles of the scattering by a spherical effective medium given by Eq. (49). The interpretation is that these resonances are very robust and depend little on the precise geometric shape of the obstacle or the strength of its potential. Therefore, to get a general idea of these resonances, it is sufficient to study the scattering by a hard sphere.
As a reminder, the hard-sphere scattering model is defined by the singular potential [57, 58, 59, 60]
$U(\boldsymbol{\mathrm{r}})=\begin{cases}+\infty&\textrm{for}~{}\left\|\boldsymbol{\mathrm{r}}\right\|\leq R\>,\\ 0&\textrm{otherwise}\>,\end{cases}$ (119)
in the wave equation (2). According to partial wave theory and in particular Eq. (117), the scattering matrix element $S_{\ell}(k)$ corresponding to Eq. (119) reads
$S_{\ell}(k)=-\frac{h^{-(d)}_{\ell}(kR)}{h^{+(d)}_{\ell}(kR)}\>.$ (120)
and thus reaches a pole when
$h^{+(d)}_{\ell}(kR)=0\>.$ (121)
Using the definition (116), Eq. (121) can directly be expressed in terms of the Hankel function
$H^{+}_{\nu}(z)=0\>,$ (122)
where $z=kR$ and the order $\nu$ is given by Eq. (50). Equation (122) means that the resonance poles of the hard sphere model are just the zeros of the Hankel function $H^{+}_{\nu}(z)$ for $z\in\mathbb{C}$. These zeros have been little considered in the literature. The NIST Handbook of Mathematical Functions [55] does not explicitly give quantitative estimates of these zeros, but refers to the more in-depth articles [85, 86]. In particular, the most successful theory describing these zeros seems to be in Ref. [86]. This is thus the main reference of this appendix.
The first zeros of the Hankel function are shown in Fig. 11.
Figure 11: First complex zeros of the Hankel function $H^{+}_{\nu}(z)$ for $\arg z\in[-\pi,0]$ and for integer and half-integer orders. The dotted curves represent the exact trajectory of the zeros when $\nu$ is continuously increased. The colored curves are given by the approximation (134) for continuous parameter $\tau\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$ and fixed values of $\nu$. The other zeros around $\arg z=-\pi$ for $\nu\in\mathbb{Z}$ are not shown.
Given the relation (50), integer orders correspond to even dimensions ($d\in\{2,4,6,\ldots\}$), and half-integer orders to odd dimensions ($d\in\{1,3,5,\ldots\}$). When $\nu$ is relatively large compared to 11 11, the zeros draw a kind of elliptic arc in the lower half-plane ($\arg z\in[-\pi,0]$). The size of this arc linearly increases with $\nu$. When $\nu$ is half an integer, the zeros are symmetrically placed with respect to the imaginary $z$ axis. It should be noted that an infinite string of zeros may also exist along the negative real axis ($\arg z=-\pi$) when the order $\nu$ is not half an integer. However, we do not consider these additional zeros in this work, because they are too far from the real positive axis ($\arg z=0$) and not relevant for the numerical simulations. These zeros are thus not shown in Fig. 11. In this section, we restrict ourselves to the arc of zeros for $\arg z\in[-\pi,0]$.
The idea to solve Eq. (122) for $z$ is to approximate it by asymptotic expressions valid for large positive $\nu$. The issue is that there is no direct asymptotic expansion for $H^{+}_{\nu}(z)$ in the lower half-plane ($\arg z\in[-\pi,0]$) reliable enough to extract zeros. This is due to the exponentially increasing behavior of $H^{+}_{\nu}(z)$ in this region, which hides the zeros. To address this issue, the method proposed in Ref. [86] is to use the analytic continuation formula for $\arg z\in[-\pi,0]$ [55]
$H^{+}_{\nu}(z)=\frac{2}{\pi}\mathrm{i}^{-\nu-1}\left[\mathrm{i}\pi I_{\nu}(\mathop{}\!\mathrm{e}^{\mathrm{i}\frac{\pi}{2}}z)+\mathop{}\!\mathrm{e}^{\mathrm{i}\pi\nu}K_{\nu}(\mathop{}\!\mathrm{e}^{\mathrm{i}\frac{\pi}{2}}z)\right]\>,$ (123)
where $I_{\nu}(z)$ and $K_{\nu}(z)$ are the modified Bessel functions [55]. The branch cut in the right-hand side of Eq. (123) is located on the semiaxis $\arg z=-\tfrac{3\pi}{2}$. The advantage of the continuation (123) is that, in the region of the zeros ($\arg z\in[-\pi,0]$), the Bessel functions are now evaluated in the right half-plane ($\left|\arg(\mathrm{i}z)\right|\leq\tfrac{\pi}{2}$) where useful asymptotic approximations for large $\nu$ can be found. In this regard, we will use the approximations
$\begin{cases}I_{\nu}(z)\xrightarrow{\nu\gg\left|z\right|}\frac{(z/2)^{\nu}}{\Gamma(\nu+1)}\exp\!\left(\frac{(z/2)^{2}}{\nu+1}\right)\>,\\ K_{\nu}(z)\xrightarrow{\nu\gg\left|z\right|}\frac{\Gamma(\nu)}{2(z/2)^{\nu}}\exp\!\left(-\frac{(z/2)^{2}}{\nu-1}\right)\>.\end{cases}$ (124)
As pointed in Ref. [86], it seems that, while being accurate enough to get two decimal places on the sought zeros, expressions (124) are not mentioned in mathematical handbooks such as Ref. [55]. Using expression (123), Eq. (122) becomes
$\ln\left(\frac{I_{\nu}(\mathrm{i}z)}{K_{\nu}(\mathrm{i}z)}\right)=\mathrm{i}\pi(\nu+\tfrac{1}{2})-\ln\pi+2\pi\mathrm{i}n\quad\forall n\in\mathbb{Z}\>,$ (125)
where we have taken the logarithm, hence the appearance of an arbitrary integer index $n$. Now, if we substitute the Bessel functions in Eq. (125) by their approximations (124), we can write
$\begin{split}&2\nu\ln(\mathrm{i}z/2)+\frac{2\nu(\mathrm{i}z/2)^{2}}{\nu^{2}-1}\\ &=\mathrm{i}\pi(\nu+\tfrac{1}{2})+\ln[\Gamma(\nu)\Gamma(\nu+1)]-\ln(2\pi)+2\pi\mathrm{i}n\>.\end{split}$ (126)
It is worth noting that Eq. (126) can already be solved exactly for $z$ by means of the Lambert function $W(z)$ defined by [55]
$W(z)\mathop{}\!\mathrm{e}^{W(z)}=z\>.$ (127)
However, further approximations can be done without loss of precision. In particular, the gamma functions can be approximated by the first three terms of the following Stirling expansion
$\begin{split}\ln[\Gamma(\nu)\Gamma(\nu+1)]&=2\nu\ln(\nu)-2\nu+\ln(2\pi)\\ &+\tfrac{1}{6\nu}+\operatorname{\mathcal{O}}(\tfrac{1}{\nu^{3}})\>.\end{split}$ (128)
Therefore, Eq. (126) becomes
$\ln\left(\frac{\mathrm{i}z}{2\nu}\right)+\frac{(\mathrm{i}z/2)^{2}}{\nu^{2}-1}=\mathrm{i}\pi\frac{\nu+\frac{1}{2}+2n}{2\nu}-1\>.$ (129)
If, in addition, we assume that $\nu$ is large enough so that $\nu^{2}-1\simeq\nu^{2}$, then it is even more appropriate to define the change of variable
$x=\frac{\mathrm{i}z}{2\nu}\>.$ (130)
In contrast to $z$ which increases linearly with $\nu$ near the zeros, the new variable $x$ remains more or less constant in absolute value. In addition, $x$ belongs to the right half-plane ($\left|\arg x\right|\leq\tfrac{\pi}{2}$). In order to simplify Eq. (129), it is convenient to define the angular parameter $\tau$ from its right-hand side
$\tau=\pi\frac{\nu+\frac{1}{2}+2n}{2\nu}\quad\forall n\in\{-\left\lfloor\nu+\tfrac{1}{4}\right\rfloor,\ldots,-1\}\>.$ (131)
The integer values of $n$ are chosen such that $\tau$ belongs to the interval $[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$. The number of zeros described by this approximation is thus $\left\lfloor\nu+\tfrac{1}{4}\right\rfloor$. However, note that the total number of zeros is not finite when $\nu$ is not half an integer. With the two notations $x$ and $\tau$, Eq. (129) becomes much more readable
$\ln(x)+x^{2}=\mathrm{i}\tau-1\>,$ (132)
where $x$ is the unknown. A first rough estimate of the solution of Eq. (132) is $x\approx\mathop{}\!\mathrm{e}^{\mathrm{i}\tau-1}$. This estimate shows that, when $n$ varies, the zeros draw an elliptic curve close to an arc of circle of radius $\mathop{}\!\mathrm{e}^{-1}$ and angles $\tau\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$ in the complex plane of $x$. In fact, it is possible to solve exactly Eq. (132) for $x$ with the Lambert $W(z)$ function. We obtain
$x=\sqrt{\frac{1}{2}W(2\mathop{}\!\mathrm{e}^{2\mathrm{i}\tau-2})}\qquad\text{for}~{}\arg x\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]\>.$ (133)
Only the principal branch of $W(z)$ must be considered here since the argument is small enough in absolute value compared to the radius of convergence of the Maclaurin series of the Lambert function, which is $\mathop{}\!\mathrm{e}^{-1}$. This is true because indeed $\tau\in\mathbb{R}$ and $2\mathop{}\!\mathrm{e}^{-2}<\mathop{}\!\mathrm{e}^{-1}$. Finally, using Eq. (130), we obtain the approximate positions of the zeros of $H^{+}_{\nu}(z)$
$z=-2\mathrm{i}\nu\sqrt{\frac{1}{2}W(2\mathop{}\!\mathrm{e}^{2\mathrm{i}\tau-2})}\>,$ (134)
where $\tau$ is given by (131). Note that other approximations not relying on the Lambert function were proposed in Ref. [86], but we do not find them useful here because, from a numerical perspective, the Lambert function can be computed very quickly in Maclaurin series for small arguments, or more generally with the Newton method [55]. Approximation (134) has proven very useful in practice to feed the numerical root-finding method applied to solve the resonance equation (49).
Appendix D Normalization of the density matrix
The purpose of this appendix is to verify that the normalization of the density matrix (64) is conserved in the time domain. Of course, this must be the case because the wave packet does not undergo absorption in our model. The purpose of the following calculation is twofold: Proving that the total probability is conserved after time $t=0$ to check the consistency of this formalism, but also finding the normalization constant. The total probability in the density matrix $\operatorname{Tr}\hat{\rho}(s)$ is given by its trace. Thus, using Eq. (64), we need to find
$\operatorname{Tr}\hat{\rho}(s)=\operatorname{Tr}\left[\hat{Q}(\omega_{+})\hat{\rho}_{0}\hat{Q}(\omega_{-})\right]\>,$ (135)
where $\omega_{\pm}=\omega_{0}\pm\tfrac{\mathrm{i}}{2}s$. The resolvents in Eq. (135) can be put next to each other in virtue of the cyclic property of trace $\operatorname{Tr}(\hat{A}\hat{B}\hat{C})=\operatorname{Tr}(\hat{B}\hat{C}\hat{A})$, which holds for trace-class operators. Then, their product is decomposed using the first resolvent identity [87]
$\hat{Q}(\omega_{-})\hat{Q}(\omega_{+})=-\frac{\hat{Q}(\omega_{+})-\hat{Q}(\omega_{-})}{\omega_{+}-\omega_{-}}\>.$ (136)
Since the Laplace variable $s$ is supposed to be much smaller than the central frequency of the wave packet ($s\ll\tfrac{2\pi}{\Delta t}\ll\omega_{0}$), it can be treated as an infinitesimal parameter in the difference $\hat{Q}(\omega_{+})-\hat{Q}(\omega_{-})$, so that Eq. (136) gets simplified in
$\hat{Q}(\omega_{-})\hat{Q}(\omega_{+})\xrightarrow{s\ll\omega_{0}}\frac{2\pi}{s}\delta(\omega_{0}-\hat{H})\>,$ (137)
where $\hat{H}$ denotes the Hamiltonian of the full system, possibly with a random potential. Therefore, Eq. (135) becomes
$\operatorname{Tr}\hat{\rho}(s)=\frac{2\pi}{s}\underbrace{\operatorname{Tr}\left[\delta(\omega_{0}-\hat{H})\hat{\rho}_{0}\right]}_{\mu(\omega_{0}\mid\hat{\rho}_{0})}\>,$ (138)
We identify the remaining trace in the right-hand side of Eq. (138) as the local density of states $\mu(\omega_{0}\mid\hat{\rho}_{0})$ in the initial state $\hat{\rho}_{0}$. This factor is just a time independent constant because it does not depend on $s$. Finally, the time evolution of the trace of the density matrix can be obtained through inverse Laplace transform
$\begin{split}\operatorname{Tr}\hat{\rho}(t)&=\frac{1}{2\pi\mathrm{i}}\int_{-\mathrm{i}\infty}^{\mathrm{i}\infty}\mathop{}\!\mathrm{e}^{st}\frac{2\pi}{s}\mu(\omega_{0}\mid\hat{\rho}_{0})\mathop{}\!\mathrm{d}s\\ &=2\pi\mu(\omega_{0}\mid\hat{\rho}_{0})\Theta(t)\>,\end{split}$ (139)
where $\Theta(t)$ denotes the Heaviside step function:
$\Theta(t)=\begin{cases}0&t<0\>,\\ 1&t\geq 0\>.\end{cases}$ (140)
The result (139) confirms that the probability is indeed conserved since the initial time $t=0$, as expected. It also gives us the normalization constant of the density matrix (64), which is just $2\pi\mu(\omega_{0}\mid\hat{\rho}_{0})$ [25].
Appendix E Propagation in free space
In this appendix, we check the consistency of the density matrix (66) in the special case of a freely propagating wave packet initially located at the position $\boldsymbol{\mathrm{r}}_{0}$ at time $t=0$. In this case, the Green function reduces to the free-space Green function (10). Thus, we have
$\left\langle\boldsymbol{\mathrm{r}}\right|\hat{\rho}(s)\left|\boldsymbol{\mathrm{r}}^{\prime}\right\rangle=G^{+}(k_{+},r){G^{+}(k_{+},r^{\prime})}^{*}\>,$ (141)
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0}\right\|$ and $r^{\prime}=\left\|\boldsymbol{\mathrm{r}}^{\prime}-\boldsymbol{\mathrm{r}}_{0}\right\|$, and the wavenumber is $k_{+}=k_{0}+\mathrm{i}\tfrac{s}{2v}$ according to Eq. (67). At large enough distance, the Green function can be approached by the asymptotic behavior (11). Therefore, the evanescent effect of the imaginary part of $k_{+}$ can be factored out as
$\left\langle\boldsymbol{\mathrm{r}}\right|\hat{\rho}(s)\left|\boldsymbol{\mathrm{r}}^{\prime}\right\rangle\simeq G^{+}(k_{0},r){G^{+}(k_{0},r^{\prime})}^{*}\mathop{}\!\mathrm{e}^{-\frac{s}{2v}(r+r^{\prime})}\>.$ (142)
The time dependence can then be restored by inverse Laplace transform, turning the decreasing exponential into a Dirac delta
$\left\langle\boldsymbol{\mathrm{r}}\right|\hat{\rho}(t)\left|\boldsymbol{\mathrm{r}}^{\prime}\right\rangle=G^{+}(k_{0},r){G^{+}(k_{0},r^{\prime})}^{*}\delta(t-\tfrac{r+r^{\prime}}{2v})\>.$ (143)
In particular, the probability density can be found on the diagonal of the density matrix, that is to say
$\rho(\boldsymbol{\mathrm{r}},t)=\left|G^{+}(k_{0},r)\right|^{2}\delta(t-\tfrac{r}{v})\>.$ (144)
Finally, the Green function in Eq. (144) can be eliminated using Eq. (11) and $v=2k_{0}$. One finds [25]
$\rho(\boldsymbol{\mathrm{r}},t)=\frac{2\pi\mu_{0}(\omega_{0}\mid\boldsymbol{\mathrm{r}}_{0})}{S_{d}r^{d-1}}\delta(r-vt)\>,$ (145)
where $\mu_{0}(\omega_{0}\mid\boldsymbol{\mathrm{r}}_{0})$ is the free-space local density of states given by Eq. (13) with $k_{0}=\sqrt{\omega_{0}}$.
The result (145) can be interpreted as a spherical distribution propagating in all directions at speed $v$. In the absence of collision with the scatterers, the free motion of the wave packet continues forever from time $t=0$. The conservation of probability is ensured by the factor $S_{d}r^{d-1}$ which is just the surface of the spherical distribution. Equation (145) is also consistent with the normalization factor $2\pi\mu(\omega_{0}\mid\hat{\rho}_{0})$ previously found in Eq. (139), hence confirming the validity of our calculation.
Appendix F Derivation of the Bethe-Salpeter equation
The integral transport equation (68) is derived in this appendix using the diagrammatic approach. This equation has the same structure as the Bethe-Salpeter equation [88] and is thus often referred to as such. For this purpose, let us look closer at the density matrix $\hat{\rho}(s)$ from Eq. (64) when we substitute the propagator $\hat{Q}(k)$ for the multiple-scattering expansion (104)
$\begin{split}&\hat{Q}(k)\hat{\rho}_{0}\hat{Q}^{\dagger}(k)\\ &=\sum_{p=0}^{\infty}\sum_{i_{1},\ldots,i_{p}}\sum_{q=0}^{\infty}\sum_{i^{\prime}_{1},\ldots,i^{\prime}_{q}}\hat{G}\hat{F}_{i_{p}}\hat{G}\cdots\hat{G}\hat{F}_{i_{1}}\hat{G}\hat{\rho}_{0}\\ &\hskip 100.00015pt\times\hat{G}^{\dagger}\hat{F}^{\dagger}_{i^{\prime}_{1}}\hat{G}^{\dagger}\cdots\hat{G}^{\dagger}\hat{F}^{\dagger}_{i^{\prime}_{q}}\hat{G}^{\dagger}\>.\end{split}$ (146)
Note that the wavenumber is implicitly complex $k=k_{+}=k_{0}+\tfrac{\mathrm{i}}{2}\gamma$ with $\gamma=s/v$. Note also that the sums over the paths $\mathcal{P}=(i_{1},\ldots,i_{p})$ and $\mathcal{Q}=(i^{\prime}_{1},\ldots,i^{\prime}_{q})$ in Eq. (146) are completely independent of each other, meaning that the indices are allowed to coincide. To remain consistent with the approximations we made in Sec. III.1, we assume that the number of scatterers is large enough so that the Twersky approximation holds. In other words, we assume that the paths $\mathcal{P}$ and $\mathcal{Q}$ involved in Eq. (146) are both self-avoiding. However, this does not prevent some indices of $\mathcal{P}$ to be equal to indices of $\mathcal{Q}$.
The diagrams of some low-order terms of Eq. (146) are shown in Fig. 12 using the notations of Fig. 9(a).
Figure 12: Expansion of the density matrix $\hat{\rho}(s)$ from Eq. (146) under the Twersky approximation using the notations of Fig. 9(a). The black dots represent the initial state $\hat{\rho}_{0}$. The list of diagrams is not exhaustive. The sums implicitly assume that the indices $i_{1},i_{2},\ldots$ are different from each other.
To make things easier to read, it is more convenient to fold the diagrams in half around the initial state $\hat{\rho}_{0}$, as done in the modern literature [25, 26, 27, 52, *Doicu2019a2, *Doicu2019a3]. The initial state $\hat{\rho}_{0}$ itself is represented by a black dot on the right side. In this way, the sequence of collisions should be chronologically read from right to left, as it was the case in Fig. 9(b). As one can see, the diagrams with links in the upper branch or in the lower branch have already been excluded from Fig. 12 according to the Twersky approximation, such as in Fig. 9(b). The only allowed diagrams with links connect the upper and lower branches. Note also that one scattering site cannot have more than one link again because of the Twersky approximation. Indeed, that would mean that one would have internal links in the other branch.
With all these rules in mind, we are able to expand the configurational average of Eq. (146) further. Let us first consider the diagrams with no link shown in the first three lines of Fig. 12. Since then all the scatterers are assumed to be different, the configurational average applies separately to each branch, and we can write
$\begin{split}\left\langle\hat{Q}(k)\hat{\rho}_{0}\hat{Q}^{\dagger}(k)\right\rangle_{0}&=\left\langle\sum_{p=0}^{\infty}\sum_{i_{1},\ldots,i_{p}}\hat{G}\hat{F}_{i_{p}}\hat{G}\cdots\hat{G}\hat{F}_{i_{1}}\hat{G}\right\rangle\hat{\rho}_{0}\\ \times&\left\langle\sum_{q=0}^{\infty}\sum_{i^{\prime}_{1},\ldots,i^{\prime}_{q}}\hat{G}^{\dagger}\hat{F}^{\dagger}_{i^{\prime}_{1}}\hat{G}^{\dagger}\cdots\hat{G}^{\dagger}\hat{F}^{\dagger}_{i^{\prime}_{q}}\hat{G}^{\dagger}\right\rangle\>.\end{split}$ (147)
The zero subscript of $\langle\cdot\rangle_{0}$ refers to the absence of link between the two branches. Note that the sums in Eq. (147) implicitly assume that no index of the first path $\mathcal{P}$ can be equal to any of the indices of the second path $\mathcal{Q}$. Now, we notice that each of the two branches in Eq. (147) is given by the average of the full propagator (104). Therefore, Eq. (147) is equivalent to
$\left\langle\hat{Q}(k)\hat{\rho}_{0}\hat{Q}^{\dagger}(k)\right\rangle_{0}=\left\langle\hat{Q}\right\rangle\hat{\rho}_{0}\left\langle\hat{Q}^{\dagger}\right\rangle\>,$ (148)
which can be easily interpreted as the average of a product of independent random variables.
Then, we consider the diagrams with a single link, partially shown in the third line of Fig. 12. One way to treat them is to fix the scatterer involved in the link, say $\boldsymbol{\mathrm{x}}_{i}$, and to split the paths $\mathcal{P}$ and $\mathcal{Q}$ in two around this point [52, *Doicu2019a2, *Doicu2019a3, 63]. Thus, we have more explicitly
$\begin{cases}\mathcal{P}=(\mathcal{P}_{1},i,\mathcal{P}_{2})=(i_{1},i_{2},\ldots,i_{p_{1}},i,j_{1},j_{2},\ldots,j_{p_{2}})\>,\\ \mathcal{Q}=(\mathcal{Q}_{1},i,\mathcal{Q}_{2})=(i^{\prime}_{1},i^{\prime}_{2},\ldots,i^{\prime}_{q_{1}},i,j^{\prime}_{1},j^{\prime}_{2},\ldots,j^{\prime}_{q_{2}})\>.\end{cases}$ (149)
Given the Twersky approximation, no other indices than $i$ are allowed to coincide. The index $i$ itself must be unique. Using the paths (149), the terms with one link in Eq. (146) can be gathered as
$\begin{split}\left\langle\hat{Q}(k)\hat{\rho}_{0}\hat{Q}^{\dagger}(k)\right\rangle_{1}&=\left\langle\sum_{i}\left(\sum_{\mathcal{P}_{1},\mathcal{P}_{2}}\hat{G}_{\mathcal{P}_{1}}\hat{F}_{i}\hat{G}_{\mathcal{P}_{2}}\right)\hat{\rho}_{0}\right.\\ &\times\left.\left(\sum_{\mathcal{Q}_{1},\mathcal{Q}_{2}}\hat{G}^{\dagger}_{\mathcal{Q}_{2}}\hat{F}^{\dagger}_{i}\hat{G}^{\dagger}_{\mathcal{Q}_{1}}\right)\right\rangle\>,\end{split}$ (150)
where the notation $\hat{G}_{\mathcal{P}}$ stands for the product
$\hat{G}_{\mathcal{P}}=\hat{G}\hat{F}_{i_{p}}\hat{G}\cdots\hat{G}\hat{F}_{i_{2}}\hat{G}\hat{F}_{i_{1}}\hat{G}\>.$ (151)
Again, the subscript of $\langle\cdot\rangle_{1}$ in Eq. (150) recalls that the diagrams contain only one link. Since all the points in $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ are statistically independent, the configurational average applies to each of them separately. They all yield the effective propagator
$\left\langle\sum_{\mathcal{P}}\hat{G}_{\mathcal{P}}\right\rangle=\left\langle\hat{Q}\right\rangle\>,$ (152)
according to Eqs. (104) and (151). In the end, only the terms $\hat{F}_{i}$ and $\hat{F}^{\dagger}_{i}$ remain because they are obviously not independent. We can now release the previously fixed scatterer $\boldsymbol{\mathrm{x}}_{i}$, and average it in space. The result reads
$\begin{split}\left\langle\hat{Q}(k)\hat{\rho}_{0}\hat{Q}^{\dagger}(k)\right\rangle_{1}=&\int_{\mathbb{R}^{d}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}}\left\langle\hat{Q}\right\rangle\hat{\rho}_{0}\\ &\times\left\langle\hat{Q}^{\dagger}\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}}\left\langle\hat{Q}^{\dagger}\right\rangle n(\boldsymbol{\mathrm{x}})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}\>.\end{split}$ (153)
where $\hat{F}_{\boldsymbol{\mathrm{x}}}=F(k)\left|\boldsymbol{\mathrm{x}}\right\rangle\left\langle\boldsymbol{\mathrm{x}}\right|$.
From these calculations, we can guess the general procedure to determine the contribution from diagrams with more links. This can be summarized in three steps.
  • First, fixing the scatterers involved in the links,
  • then averaging over the independent paths, leading to factors $\langle\hat{Q}\rangle$,
  • and, finally, averaging over the previously fixed scatterers.
In particular, the second step is illustrated in Fig. 13.
Figure 13: Resummation of the diagrams with no link for the average density matrix $\langle\hat{\rho}(s)\rangle$. (a) Individual resummation of diagrams with fixed number of vertices in the lower arm. (b) Final resummation of all the contributions from panel (a).
In fact, between any fixed scatterers, we may have an arbitrary number of collisions, which, on average, leads to the effective propagator $\langle\hat{Q}\rangle$. Therefore, everything is going as if $\hat{G}$ is simply replaced by $\langle\hat{Q}\rangle$ upon averaging. Finally, only links remain visible. For instance, the diagrams with two links, including those in the last two lines of Fig. 12, can also be summed up in this way. However, we must be careful that the links may cross each other. This leads to two separate kinds of terms whose resummation reads
$\begin{split}\left\langle\hat{\rho}\right\rangle_{2}&=\iint_{\mathbb{R}^{2d}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}_{2}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}_{1}}\left\langle\hat{Q}\right\rangle\hat{\rho}_{0}\\ &\times\left\langle\hat{Q}^{\dagger}\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}_{1}}\left\langle\hat{Q}^{\dagger}\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}_{2}}\left\langle\hat{Q}^{\dagger}\right\rangle n(\boldsymbol{\mathrm{x}}_{1})n(\boldsymbol{\mathrm{x}}_{2})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{1}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{2}\\ &+\iint_{\mathbb{R}^{2d}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}_{2}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}_{1}}\left\langle\hat{Q}\right\rangle\hat{\rho}_{0}\\ &\times\left\langle\hat{Q}^{\dagger}\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}_{2}}\left\langle\hat{Q}^{\dagger}\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}_{1}}\left\langle\hat{Q}^{\dagger}\right\rangle n(\boldsymbol{\mathrm{x}}_{1})n(\boldsymbol{\mathrm{x}}_{2})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{1}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}_{2}\>.\end{split}$ (154)
Note the order of indices of $\hat{F}_{\boldsymbol{\mathrm{x}}_{1}}$ and $\hat{F}_{\boldsymbol{\mathrm{x}}_{2}}$ in these two terms. The new expansion of $\left\langle\hat{\rho}(s)\right\rangle$ is graphically represented in Fig. 14(b) using the symbol in Fig. 14(a). The upper and lower legs of the symbol in Fig. 14(a) depicts the left and right inputs of the operator $\left\langle\hat{\rho}(s)\right\rangle$, respectively.
Figure 14: (a) Diagram for the average density matrix. (b) Full expansion of $\langle\hat{\rho}(s)\rangle$. (c) Expansion of $\langle\hat{\rho}(s)\rangle$ in the ladder approximation. (d) Recursive form of $\langle\hat{\rho}(s)\rangle$ in the ladder approximation, also referred to as the Bethe-Salpeter equation.
As one can see, Fig. 14(b) is very similar to Fig. 12, except that the unlinked scatterers are omitted and replaced by the effective propagator $\langle\hat{Q}\rangle$ or $\langle\hat{Q}^{\dagger}\rangle$. This is a consequence of the resummation illustrated in Fig. 13.
As already pointed out, some of the diagrams in Fig. 14(b) possess cross links, especially in the last two lines. We will simply neglect them in our calculation. This precise step is known as the ladder approximation or the diffuson approximation [^3] This terminology must not be confused with the diffusion approximation presented in Sec. IV.3 (note the “i”). in the literature [24, 25, 26, 69]. The former name comes from the structure of the diagrams shown in Fig. 14(c). The validity of this approximation is difficult to justify rigorously, but it seems relevant to describe the wave propagation in the weak scattering regime (69) in the absence of localization [49, *Fante1982, 89, 25]. In fact, the ladder diagrams in Fig. 14(c) account for the incoherent propagation of the wave, while the cross diagrams in the end of Fig. 14(b) account for the coherent effects, such as the coherent backscattering [89, 90, 91, 92, 23, 25, 27, 69, 7]. As explained near Tab. 1, it is reasonable to neglect these coherent effects in gaseous media.
Finally, the ladder diagrams in Fig. 14(c) can be summed up using the same kind of recursive equation as Eq. (109). In this way, one obtains the following integral equation of the wave transport [28, 68, 19, 20, 21, 24, 25, 26, 69]
$\begin{split}\left\langle\hat{\rho}(s)\right\rangle&=\left\langle\hat{Q}\right\rangle\hat{\rho}_{0}\left\langle\hat{Q}^{\dagger}\right\rangle\\ &+\int_{\mathbb{R}^{d}}\left\langle\hat{Q}\right\rangle\hat{F}_{\boldsymbol{\mathrm{x}}}\left\langle\hat{\rho}(s)\right\rangle\hat{F}^{\dagger}_{\boldsymbol{\mathrm{x}}}\left\langle\hat{Q}^{\dagger}\right\rangle n(\boldsymbol{\mathrm{x}})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}\>.\end{split}$ (155)
This equation is very similar to the Bethe-Salpeter equation used in relativistic quantum field theory [88], and is often referred to as such, especially in the radiative transfer literature [49, *Fante1982, 89, 52, *Doicu2019a2, *Doicu2019a3]. In fact, Eq. (155) was first derived in Foldy’s seminal paper on the model of point scatterers [28, 29], but his approximations remained vaguely justified until the thorough work by Twersky [63] previously mentioned in Appendix A.
Equation (155) is graphically illustrated by the diagram of Fig. 14(d). We briefly recall that the right-hand side of Eq. (155) depends on $s$ through the fact that the disorder-averaged Green operators $\langle\hat{Q}\rangle$ are implicitly evaluated at the complex wavenumber $k_{+}=k_{0}+\tfrac{\mathrm{i}}{2}\gamma$, as discussed in Sec. IV.1.
At the end of the day, we ultimately obtain Eq. (68) by projecting Eq. (155) on the position basis with $\left\langle\boldsymbol{\mathrm{r}}\right|\cdots\left|\boldsymbol{\mathrm{r}}^{\prime}\right\rangle$ and by assuming the initial position state $\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|$.
Appendix G Relationship with the linear Boltzmann equation
The integral transport equation (72) is here shown to be consistent with the famous linear Boltzmann equation [34, 33, 25]
$\begin{split}&\frac{\partial f}{\partial t}+v\boldsymbol{\mathrm{\Omega}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}f\\ &=nv\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega^{\prime}\frac{\mathop{}\!\mathrm{d}\sigma}{\mathop{}\!\mathrm{d}\Omega}(\boldsymbol{\mathrm{\Omega}}\mid\boldsymbol{\mathrm{\Omega}}^{\prime})\big{(}f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}}^{\prime},t)-f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)\big{)}\>,\end{split}$ (156)
where $\textstyle\frac{\mathop{}\!\mathrm{d}\sigma}{\mathop{}\!\mathrm{d}\Omega}$ is the differential cross section of a point scatterer, and $f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)$ is the distribution function of a classical particle in the phase space, $\boldsymbol{\mathrm{r}}$ being the position, and $\boldsymbol{\mathrm{\Omega}}$ the direction of motion. The function $f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)$ is closely related to the Wigner function of the wavefunction $\psi(\boldsymbol{\mathrm{r}},t)$ averaged over the random configurations of the disorder.
The loss term in the right-hand side of Eq. (156) can be exactly evaluated, since $f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)$ does not depend on the integrated direction $\boldsymbol{\mathrm{\Omega}}^{\prime}$. This integral lets appear the total cross section $\sigma$. Putting this term in the left-hand side, we get
$\begin{split}&\left(\frac{\partial}{\partial t}+v\boldsymbol{\mathrm{\Omega}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}+n\sigma v\right)f\\ &=nv\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega^{\prime}\frac{\mathop{}\!\mathrm{d}\sigma}{\mathop{}\!\mathrm{d}\Omega}(\boldsymbol{\mathrm{\Omega}}\mid\boldsymbol{\mathrm{\Omega}}^{\prime})f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}}^{\prime},t)\>.\end{split}$ (157)
We want to derive an integral transport equation which does not depend anymore on the particle direction of motion $\boldsymbol{\mathrm{\Omega}}$. To this end, we will use the Green function method. We interpret the brackets in the left-hand side of Eq. (157) as a linear operator acting on $f$ to be inverted. Thus, we consider the integral kernel given by the equation
$\begin{split}&\left(\frac{\partial}{\partial t}+v\boldsymbol{\mathrm{\Omega}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}+n\sigma v\right)K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},\boldsymbol{\mathrm{\Omega}}^{\prime},t^{\prime})\\ &=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})\delta(\boldsymbol{\mathrm{\Omega}}-\boldsymbol{\mathrm{\Omega}}^{\prime})\delta(t-t^{\prime})\>.\end{split}$ (158)
The function $K$ is generally referred to as the transport kernel [34]. The solution of Eq. (158) reads
$\begin{split}&K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},\boldsymbol{\mathrm{\Omega}}^{\prime},t^{\prime})\\ &=\mathop{}\!\mathrm{e}^{-n\sigma\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}-v\boldsymbol{\mathrm{\Omega}}(t-t^{\prime}))\delta(\boldsymbol{\mathrm{\Omega}}-\boldsymbol{\mathrm{\Omega}}^{\prime})\>.\end{split}$ (159)
Note that this kernel is supposed to be identically zero for $t<t^{\prime}$. This can be realized with a Heaviside step function $\Theta(t-t^{\prime})$, but we leave it implicit.
The function (159) can be interpreted as the distribution of a particle located at $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{r}}^{\prime}+v\boldsymbol{\mathrm{\Omega}}(t-t^{\prime})$ for $t\geq t^{\prime}$ in position space and $\boldsymbol{\mathrm{\Omega}}=\boldsymbol{\mathrm{\Omega}}^{\prime}$ in direction space. This is the probability distribution that the particle does not undergo any collision after distance $\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$ from the previous collision at $\boldsymbol{\mathrm{r}}^{\prime}$ [25]. The exponential prefactor comes from the constant collision rate $n\sigma v$ in the left-hand side of Eq. (158). Since the collisions constitute a Poisson process and are less and less unavoidable over time, this probability exponentially vanishes. In some way, this decay can be understood as the absorption of the particle in a collision event. The particle is then recreated by the gain term in the right-hand side of Eq. (158). Of course, this is a strictly mathematical interpretation, and it does not mean that the model assumes any absorption effects by the gas.
Furthermore, we also define in advance the following integral kernel averaged over the initial direction of the particle
$\begin{split}K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})&=\frac{1}{S_{d}}\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega^{\prime}K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},\boldsymbol{\mathrm{\Omega}}^{\prime},t^{\prime})\\ &=\frac{1}{S_{d}}\mathop{}\!\mathrm{e}^{-n\sigma\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}-v\boldsymbol{\mathrm{\Omega}}(t-t^{\prime}))\>,\end{split}$ (160)
where $S_{d}$ is the surface of the unit $d$-ball given in Eq. (1). This kernel describes an initially spherical wave packet, but still depends on the direction $\boldsymbol{\mathrm{\Omega}}$. If we assume that the initial state is such a spherical state
$f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t_{0})=\frac{1}{S_{d}}\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})\>,$ (161)
then the Boltzmann equation (157) can be rewritten in integral form [34]
$\begin{split}f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)&=K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}_{0},t_{0})\\ &+\int_{t_{0}}^{t}\mathop{}\!\mathrm{d}t^{\prime}\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega^{\prime}K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},\boldsymbol{\mathrm{\Omega}}^{\prime},t^{\prime})\\ &\qquad\times nv\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega^{\prime\prime}\frac{\mathop{}\!\mathrm{d}\sigma}{\mathop{}\!\mathrm{d}\Omega}(\boldsymbol{\mathrm{\Omega}}^{\prime}\mid\boldsymbol{\mathrm{\Omega}}^{\prime\prime})f(\boldsymbol{\mathrm{r}}^{\prime},\boldsymbol{\mathrm{\Omega}}^{\prime\prime},t^{\prime})\>.\end{split}$ (162)
In order to match the point scattering model of Sec. II.1, we assume that the cross section is isotropic, i.e., $\tfrac{\mathop{}\!\mathrm{d}\sigma}{\mathop{}\!\mathrm{d}\Omega}=\tfrac{\sigma}{S_{d}}$. The integral over $\boldsymbol{\mathrm{\Omega}}^{\prime\prime}$ in Eq. (162) thus becomes trivial and leads to the particle density
$\rho(\boldsymbol{\mathrm{r}},t)=\oint_{\mathcal{S}_{d}}f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)\mathop{}\!\mathrm{d}\Omega\>.$ (163)
In addition, the outer integral over $\boldsymbol{\mathrm{\Omega}}^{\prime}$ can be simplified using the direction-averaged kernel (160). We get
$\begin{split}f(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t)&=K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}_{0},t_{0})\\ &+\int_{t_{0}}^{t}\mathop{}\!\mathrm{d}t^{\prime}\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})n\sigma v\rho(\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})\>.\end{split}$ (164)
At this point, we can elegantly eliminate the dependency on the direction of motion in Eq. (164) by integration over $\boldsymbol{\mathrm{\Omega}}$ in both members. The new direction-independent kernel is in this regard
$K(\boldsymbol{\mathrm{r}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})=\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega\>K(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{\Omega}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})\>.$ (165)
The integral (165) can be evaluated from Eq. (160) with the following property of the Dirac delta
$\oint_{\mathcal{S}_{d}}\mathop{}\!\mathrm{d}\Omega\>\delta^{(d)}(\boldsymbol{\mathrm{r}}-vt\boldsymbol{\mathrm{\Omega}})=\frac{\delta(\left\|\boldsymbol{\mathrm{r}}\right\|-vt)}{\left\|\boldsymbol{\mathrm{r}}\right\|^{d-1}}\>.$ (166)
The result is
$K(\boldsymbol{\mathrm{r}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})=\frac{\mathop{}\!\mathrm{e}^{-n\sigma\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\delta(\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|-v(t-t^{\prime}))\>,$ (167)
and really represents a spherical shell expanding in every direction at constant speed $v$. The integral equation (164) becomes
$\begin{split}\rho(\boldsymbol{\mathrm{r}},t)&=K(\boldsymbol{\mathrm{r}},t\mid\boldsymbol{\mathrm{r}}_{0},t_{0})\\ &+\int_{t_{0}}^{t}\mathop{}\!\mathrm{d}t^{\prime}\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}K(\boldsymbol{\mathrm{r}},t\mid\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})n\sigma v\rho(\boldsymbol{\mathrm{r}}^{\prime},t^{\prime})\>,\end{split}$ (168)
which no longer depends on the direction of motion. Note that the elimination of the direction dependence is fundamentally allowed by the assumption of isotropic cross section.
Finally, one notices that the transport equation (168) is stationary in the sense that the physical properties of the scattering medium do not depend on time. Consequently, it is useful to apply an integral transform, such as the Laplace transform, to Eq. (168):
$\rho(\boldsymbol{\mathrm{r}},s)=v\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\>\mathop{}\!\mathrm{e}^{-st}\rho(\boldsymbol{\mathrm{r}},t)\>.$ (169)
The factor of $v$ in Eq. (169) is motivated by dimensional reasons. The Laplace transform also applies to the transport kernel, which becomes
$\begin{split}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})&=v\int_{0}^{\infty}\mathop{}\!\mathrm{d}t\>\mathop{}\!\mathrm{e}^{-st}K(\boldsymbol{\mathrm{r}},t\mid\boldsymbol{\mathrm{r}}^{\prime},0)\\ &=\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\>,\end{split}$ (170)
where we used the notation $\gamma=\tfrac{s}{v}$. Note that we retrieve in Eq. (170) the behavior of the transport kernel (71). The advantage of the Laplace transform is that the time convolution in Eq. (168) reduces to a simple product. The result is
$\rho(\boldsymbol{\mathrm{r}},s)=K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+\int_{\mathbb{R}^{d}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n\sigma\rho(\boldsymbol{\mathrm{r}}^{\prime},s)\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (171)
which exactly coincides with the integral transport equation (72). This confirms the equivalence between equations (72) and (156).
Appendix H Coherence length at equilibrium
In this appendix, we obtain the expression of the average density matrix at equilibrium in the special case of an infinite and homogeneous medium. From this expression, we deduce the coherence length (76). First, we assume that the initial state of the particle, $\hat{\rho}_{0}$, is uniformly distributed in space
$\hat{\rho}_{0}=\rho_{0}\int_{\mathbb{R}^{d}}\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}_{0}=\rho_{0}\hat{1}\>.$ (172)
The prefactor $\rho_{0}$ in Eq. (172) is arbitrary. We will fix it at the end of the calculation. Given these assumptions, we consider the following ansatz for the average density matrix
$\left\langle\hat{\rho}(s)\right\rangle=A\left\langle\hat{Q}(k)\right\rangle\left\langle\hat{Q}^{\dagger}(k)\right\rangle\>,$ (173)
where $A$ is a normalization factor that we will also fix at the end. Note that the product of effective propagators in Eq. (173) can be expanded using the first resolvent identity [87]
$\begin{split}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle&=-\frac{\left\langle\hat{Q}(k)\right\rangle-\left\langle\hat{Q}^{\dagger}(k)\right\rangle}{2\mathrm{i}\operatorname{Im}(k^{2}-nF)}\\ &=-\frac{1}{k_{0}(\gamma+n\sigma)}\operatorname{Im}\left\langle\hat{Q}(k)\right\rangle\>,\end{split}$ (174)
where we have used $\gamma=2{k}_{\rm i}$ and $\operatorname{Im}F(k)\simeq-k_{0}\sigma(k_{0})$ in the weak scattering regime. Inserting the ansatz (173) into the Bethe-Salpeter equation (155), we get
$\begin{split}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle&=\frac{\rho_{0}}{A}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle+n\left|F(k)\right|^{2}\\ \times&\int_{\mathbb{R}^{d}}\left\langle\hat{Q}\right\rangle\left|\boldsymbol{\mathrm{x}}\right\rangle\left\langle\boldsymbol{\mathrm{x}}\right|\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle\left|\boldsymbol{\mathrm{x}}\right\rangle\left\langle\boldsymbol{\mathrm{x}}\right|\left\langle\hat{Q}^{\dagger}\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}\>.\end{split}$ (175)
Since the medium is infinite, the effective propagator $\langle\hat{Q}(k)\rangle$ is given by Eq. (40) and the diagonal matrix element $\left\langle\boldsymbol{\mathrm{x}}\right|\langle\hat{Q}\rangle\langle\hat{Q}^{\dagger}\rangle\left|\boldsymbol{\mathrm{x}}\right\rangle$ does not depend on the position $\boldsymbol{\mathrm{x}}$. In addition, using Eq. (174), we find in the weak scattering regime
$\left\langle\boldsymbol{\mathrm{x}}\right|\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle\left|\boldsymbol{\mathrm{x}}\right\rangle\simeq\frac{\pi\mu(k_{0})}{k_{0}(\gamma+n\sigma)}\>.$ (176)
Thus, using Eq. (176) and the resolution of identity $\textstyle\hat{1}=\int_{\mathbb{R}^{d}}\left|\boldsymbol{\mathrm{x}}\right\rangle\left\langle\boldsymbol{\mathrm{x}}\right|\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{x}}$, the integral equation (175) reduces to
$\begin{split}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle&=\frac{\rho_{0}}{A}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle\\ &+n\left|F(k)\right|^{2}\frac{\pi\mu(k_{0})}{k_{0}(\gamma+n\sigma)}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle\>.\end{split}$ (177)
Finally, we use the point cross section formula (14) to obtain
$\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle=\frac{\rho_{0}}{A}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle+\frac{n\sigma}{\gamma+n\sigma}\left\langle\hat{Q}\right\rangle\left\langle\hat{Q}^{\dagger}\right\rangle\>,$ (178)
which is indeed true if the normalization factor $A$ satisfies
$\frac{A}{\rho_{0}}=\frac{1}{1-\frac{n\sigma}{\gamma+n\sigma}}=\frac{1}{\gamma}(\gamma+n\sigma)\>.$ (179)
Therefore, combining Eqs. (174) and (179) into the density matrix of Eq. (173), we can write
$\left\langle\hat{\rho}(s)\right\rangle=-\frac{\rho_{0}}{k_{0}\gamma}\operatorname{Im}\left\langle\hat{Q}(k)\right\rangle\>.$ (180)
The result (180) still depends on time through the variable $\gamma=s/v$ which appears in the denominator but also in $k=k_{0}+\tfrac{\mathrm{i}}{2}\gamma$. Here, we are not interested in the possible transient of the density matrix, but only in its long time limit. This limit can be determined without taking the inverse Laplace transform in time thanks to the final value theorem
$\hat{\rho}_{\infty}=\lim_{t\rightarrow\infty}\left\langle\hat{\rho}(t)\right\rangle=\lim_{s\rightarrow 0^{+}}s\left\langle\hat{\rho}(s)\right\rangle\>.$ (181)
We quickly find from Eq. (180)
$\hat{\rho}_{\infty}=-\frac{\rho_{0}v}{k_{0}}\operatorname{Im}\left\langle\hat{Q}(k_{0})\right\rangle\>.$ (182)
In the following calculations, we will set the arbitrary normalization factor to $\rho_{0}=k_{0}/v$ to get rid of the irrelevant prefactors in Eq. (182). Using Eq. (40) and the far-field behavior (11), one deduces that the imaginary part of the effective Green function is essentially the same as the free Green function but with an additional exponential decay factor [25]
$\begin{split}-\operatorname{Im}\left\langle Q^{+}(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\right\rangle&=-\operatorname{Im}G^{+}(\kappa,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\\ &\simeq I(k_{0},r)\mathop{}\!\mathrm{e}^{-\operatorname{Im}\kappa r}\>,\end{split}$ (183)
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$ and $I(k,r)=-\operatorname{Im}G^{+}(k,r)$. The imaginary part of the effective wavenumber $\kappa(k)$ in Eq. (183) is equal to $\tfrac{n\sigma}{2}$ according to Eq. (43) and the fact that ${k}_{\rm i}\rightarrow 0^{+}$ due to Eq. (181). Thus, the density matrix (182) reads in position space
$\rho_{\infty}(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})=I(k_{0},r)\mathop{}\!\mathrm{e}^{-\frac{n\sigma}{2}r}\>.$ (184)
This expression is even more explicit using the far-field behavior (11). In this way, we obtain the asymptotic expression (75).
It is worth noting that $I(k_{0},r)$ in Eq. (184) corresponds to the free-space density matrix, because the same result would have been found from the uniform statistical mixture of momentum states
$\hat{\rho}=\oint_{\mathcal{S}_{d}}\left|k_{0}\boldsymbol{\mathrm{\Omega}}\right\rangle\left\langle k_{0}\boldsymbol{\mathrm{\Omega}}\right|\mathop{}\!\mathrm{d}\Omega\>.$ (185)
According to Eqs. (17) and (18) of Ref. [46], this would lead to the position-space matrix $\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\propto I(k_{0},\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|)$, which is close to Eq. (184) but without the exponential factor $\mathop{}\!\mathrm{e}^{-\frac{n\sigma}{2}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}$. Alternatively, a statistical mixture in position space could have been constructed by integrating the free Green function $G^{+}(k_{0},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ over the source point $\boldsymbol{\mathrm{r}}_{0}$. This would have given us [^4] Strictly speaking, this integral does not converge, hence the vague proportionality signs. In principle, it should be regularized using a finite integration volume or giving to the wavenumber a small positive imaginary part, but this is not shown here.
$\begin{split}\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})&\propto\int_{\mathbb{R}^{d}}G^{+}(k_{0},\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0}){G^{+}(k_{0},\boldsymbol{\mathrm{r}}^{\prime}\mid\boldsymbol{\mathrm{r}}_{0})}^{*}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}_{0}\\ &\propto I(k_{0},\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|)\>,\end{split}$ (186)
which is again very close to Eq. (184) but without the factor $\mathop{}\!\mathrm{e}^{-\frac{n\sigma}{2}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}$.
Finally, the coherence length can be explicitly obtained from definition (74) using the far-field behavior (75) of the density matrix. In the weak scattering regime (69), the fast oscillations of the sine function in Eq. (75) can be asymptotically approached by the average value $\sin^{2}x=\tfrac{1}{2}$, and the integrals involved in Eq. (74) read
$\begin{cases}\int_{\mathbb{R}^{d}}\left|\rho_{\infty}(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{0}})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}&\xrightarrow{k_{0}\gg n\sigma}\frac{\pi\mu(k_{0})}{2k_{0}n\sigma}\>,\\ \int_{\mathbb{R}^{d}}\left\|\boldsymbol{\mathrm{r}}\right\|^{2}\left|\rho_{\infty}(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{0}})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}&\xrightarrow{k_{0}\gg n\sigma}\frac{\pi\mu(k_{0})}{k_{0}(n\sigma)^{3}}\>.\end{cases}$ (187)
In Eq. (187), we have set the arbitrary position to $\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$, because anyway the state (184) is invariant under spatial translations due to the uniform initial condition (172). From Eqs. (74) and (187), we obtain the coherence length
$\Lambda(\boldsymbol{\mathrm{r}}_{0})^{2}=\frac{2}{(n\sigma)^{2}}=2\ell_{\rm s}^{2}\>,$ (188)
and thus the result (76).
Appendix I Derivation of the diffusion equation
One way to the diffusion approximation consists in estimating the convolution integral in Eq. (72)
$K*(n\sigma\left\langle\rho\right\rangle)=n\sigma\int_{\mathcal{V}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (189)
where $\mathcal{V}=\mathcal{B}_{d}(R)$ denotes the spherical volume of the medium.
Diffusion in the bulk
We first look at the bulk of the medium, that is to say, far away from the boundary ($R-\left\|\boldsymbol{\mathrm{r}}\right\|\gg\ell_{\rm s}$). In this region, it is reasonable to extend the integral to $\mathbb{R}^{d}$, hence neglecting the possible edge effects. We will consider the special case of the boundary equation later. Omitting the constant prefactor $n\sigma$, the convolution integral (189) reads
$K*\left\langle\rho\right\rangle=\int_{\mathbb{R}^{d}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (190)
On length scales much larger than the mean free path ($r\gg\ell_{\rm s}$), an excellent approximation of the kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ is
$K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\simeq\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\>,$ (191)
according to Eq. (71). Since $n\sigma$ is very large, the kernel $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$ looks like the Dirac delta function $\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime})$, but with a nonzero variance. To exploit this feature, we expand the density $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$ in the neighborhood of the observation point $\boldsymbol{\mathrm{r}}$
$\begin{split}\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle&=\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+(r^{\prime}_{i}-r_{i})\frac{\partial}{\partial r_{i}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle\\ &+\frac{1}{2!}(r^{\prime}_{i}-r_{i})(r^{\prime}_{j}-r_{j})\frac{\partial}{\partial r_{i}}\frac{\partial}{\partial r_{j}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\cdots\>,\end{split}$ (192)
where we have assumed the implicit summation of repeated indices. We hope that the first three terms of Eq. (192) will suffice to approach $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$. The convolution integral (190) then amounts to evaluating the first moments of the distribution $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})$. The zero-order moment reads
$\int_{\mathbb{R}^{d}}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=\frac{1}{\gamma+n\sigma}\>.$ (193)
This can be easily shown using the translational and rotational invariances of the integral and the fact that $\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=S_{d}(r^{\prime})^{d-1}\mathop{}\!\mathrm{d}r^{\prime}$. The remaining integral is just $\textstyle\int_{0}^{\infty}\mathop{}\!\mathrm{e}^{-\beta x}\mathop{}\!\mathrm{d}x=\frac{1}{\beta}$. The first-order moment
$\int_{\mathbb{R}^{d}}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}(\boldsymbol{\mathrm{r}}^{\prime}-\boldsymbol{\mathrm{r}})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{0}}\>,$ (194)
is equal to the zero vector, because the distribution is symmetric with respect to the point $\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{r}}$. This also means that the particle does not undergo any drift or external forces. In fact, all the odd order moments are zero for the same reason. The second-order moment, which is a $3\times 3$ matrix, reads
$\int_{\mathbb{R}^{d}}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}(r^{\prime}_{i}-r_{i})(r^{\prime}_{j}-r_{j})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=\frac{2!}{(\gamma+n\sigma)^{3}}\frac{\delta_{ij}}{d}\>.$ (195)
The off-diagonal elements ($i\neq j$) are zero because of the spherical symmetry of the distribution. In addition, the three diagonal elements ($i=j$) are equal for the same reason. These elements can be calculated in spherical coordinates, letting $r_{i}=r\cos\theta$, but we do not show it here explicitly. Combining Eqs. (193), (194) and (195), the convolution integral (190) is approached by
$K*\left\langle\rho\right\rangle\simeq\frac{\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{\gamma+n\sigma}+\frac{\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{d(\gamma+n\sigma)^{3}}\>.$ (196)
The integral transport equation (72) then becomes
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+n\sigma\left(\frac{\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{\gamma+n\sigma}+\frac{\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{d(\gamma+n\sigma)^{3}}\right)\>.$ (197)
Multiplying both sides of Eq. (197) by $(\gamma+n\sigma)$ quickly leads to Eq. (77).
Leakage boundary conditions
Aside from the equation in the bulk, in a finite medium, we also need a condition on the density at the boundary to describe the fact that the wave can freely exit the medium. Since the medium is much larger than the mean free path, it can be approached by a semi-infinite region
$\mathcal{V}\simeq\{\boldsymbol{\mathrm{r}}\in\mathbb{R}^{d}\mid\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{n}}\leq 0\}\>,$ (198)
where $\boldsymbol{\mathrm{n}}$ is the outward-pointing normal vector to the boundary of the medium. This assumption is motivated by the fact that, at the scale of the mean free path, the boundary is perceived as nearly flat. In this way, the boundary is just an infinite plane of equation $\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{n}}=0$. In order to obtain a boundary condition on $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$, we will approximate the integral transport equation (72) for any point $\boldsymbol{\mathrm{r}}$ on the boundary. More specifically, we want to estimate the integral
$K*\left\langle\rho\right\rangle=\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (199)
where the integration domain covers the semi-infinite region (198). As done in Eq. (192), we expand the density $\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle$ in Taylor series at $\boldsymbol{\mathrm{r}}^{\prime}=\boldsymbol{\mathrm{r}}$
$\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle=\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+(\boldsymbol{\mathrm{r}}^{\prime}-\boldsymbol{\mathrm{r}})\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\cdots\>.$ (200)
We neglect the higher-order terms in Eq. (200), because we do not want the boundary condition to possess a second order derivative. Indeed, the diffusion equation (197) is of second order in space. Note that, in Eq. (199), the observation point $\boldsymbol{\mathrm{r}}$ is supposed to lie on the boundary, i.e., to satisfy $\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{n}}=0$. Without loss of generality, we can set $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{0}}$ due to the translational symmetry of Eq. (199) in the plane of the boundary. The zero-order moment of the kernel reads
$\begin{split}\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}K(\boldsymbol{\mathrm{0}}\mid\boldsymbol{\mathrm{r}}^{\prime})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}&=\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)r^{\prime}}}{S_{d}(r^{\prime})^{d-1}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\\ &=\frac{1}{2(\gamma+n\sigma)}\>.\end{split}$ (201)
This is just half the result (193) since it is integrated over half the space. The first-order moment
$\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}\boldsymbol{\mathrm{r}}^{\prime}K(\boldsymbol{\mathrm{0}}\mid\boldsymbol{\mathrm{r}}^{\prime})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}\boldsymbol{\mathrm{r}}^{\prime}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)r^{\prime}}}{S_{d}(r^{\prime})^{d-1}}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (202)
requires more caution because it is not spherically symmetric. Thus, we need to resort to spherical coordinates. Using the differential volume element in spherical coordinates
$\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}=S_{d-1}(\sin\theta)^{d-2}r^{d-1}\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}r\>,$ (203)
and projecting over the normal vector $\boldsymbol{\mathrm{n}}$, Eq. (202) becomes
$\begin{split}&\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}(\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}})K(\boldsymbol{\mathrm{0}}\mid\boldsymbol{\mathrm{r}}^{\prime})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\\ &=\frac{S_{d-1}}{S_{d}}\int_{\frac{\pi}{2}}^{\pi}\mathop{}\!\mathrm{d}\theta\,\cos\theta(\sin\theta)^{d-2}\int_{0}^{\infty}\mathop{}\!\mathrm{d}r^{\prime}\,r^{\prime}\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)r^{\prime}}\>.\end{split}$ (204)
In Eq. (204), $\theta$ denotes the angle between $\boldsymbol{\mathrm{r}}^{\prime}$ and the normal $\boldsymbol{\mathrm{n}}$. The radial and angular integrals of Eq. (204) are elementary, leading to
$\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}(\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}})K(\boldsymbol{\mathrm{0}}\mid\boldsymbol{\mathrm{r}}^{\prime})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=\frac{S_{d-1}}{S_{d}}\frac{-1}{d-1}\frac{1}{(\gamma+n\sigma)^{2}}\>.$ (205)
Restoring the arbitrary direction of the normal $\boldsymbol{\mathrm{r}}$ and using the fact that $V_{d}=S_{d}/d$, we finally obtain from Eq. (205)
$\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}\boldsymbol{\mathrm{r}}^{\prime}K(\boldsymbol{\mathrm{0}}\mid\boldsymbol{\mathrm{r}}^{\prime})\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}=-\boldsymbol{\mathrm{n}}\frac{V_{d-1}}{S_{d}}\frac{1}{(\gamma+n\sigma)^{2}}\>.$ (206)
Due to the rotational symmetry of the integral in the plane of the boundary, there is no component of the first-order moment (206) parallel to the boundary. In addition, note the minus sign in the right-hand side of Eq. (206) which is a direct consequence of the integration domain $\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0$. Combining Eqs. (201) and (206), the convolution integral (199) can be approximated by
$\begin{split}&\int_{\boldsymbol{\mathrm{r}}^{\prime}\cdot\boldsymbol{\mathrm{n}}\leq 0}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\\ &\simeq\frac{\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{2(\gamma+n\sigma)}-\frac{V_{d-1}}{S_{d}}\frac{\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{(\gamma+n\sigma)^{2}}\>.\end{split}$ (207)
Now, we can substitute the approximation (207) into the integral equation (72). The source term $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ can be neglected, because we assume that it is located far away from the boundary, typically somewhere in the bulk. This gives us
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=n\sigma\left(\frac{\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{2(\gamma+n\sigma)}-\frac{V_{d-1}}{S_{d}}\frac{\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle}{(\gamma+n\sigma)^{2}}\right)\>,$ (208)
which, after some rearrangements, leads to Eq. (84).